### v1

# How to Make an Action Better

Abstract
For two actions in a decision problem, $a$ and $b$, each of which that produces a state-dependent monetary reward, we study how to robustly make action $a$ more attractive. Action $\hat{a}$ improves upon $a$ in this manner if the set of beliefs at which $a$ is preferred to $b$ is a subset of the set of beliefs at which $\hat{a}$ is preferred to $b$, irrespective of the risk-averse agent’s utility function (in money). We provide a full characterization of this relation and discuss applications in bilateral trade and insurance.
1 Introduction
Enter Alice. We know that given a choice between actions $a$ and $b$–the former of which pays out some (monetary) prize amount $\alpha_{S}$ if the weather is sunny and $\alpha_{R}$ if the weather is rainy; and the latter of which, $\beta_{S}$ and $\beta_{R}$ in sun and rain, respectively–she prefers $a$. We do not know, however, Alice’s belief about the probability of sun, nor do we know her utility function (over wealth). Given these data and facts, and under the assumption that her utility function is increasing in wealth, what are the properties of a third action $\hat{a}$, so that she must also prefer it to $b$? What if we further assume that Alice is risk-averse?
The knee-jerk response to these questions is that it is obvious: when Alice’s utility is known only to be within the class of increasing utility functions, it must be that $\hat{a}$ corresponds to a first-order stochastic dominance improvement over $a$. Likewise, when Alice is known to be risk averse, it must be that $\hat{a}$ corresponds to a second-order stochastic dominance improvement over $a$. However, these answers do not stand up to scrutiny–in particular, we do not know the lotteries produced by Alice’s subjective belief, so we cannot speak directly of dominance of lotteries.
In our two main results, Theorems 3.7 and 3.9, we fully characterize these relations in terms of the primitives of the environment–the state-dependent payoffs to actions $a$, $\hat{a}$, and $b$. This characterization is distribution-free and almost ordinal–the lone cardinality specification needed merely concerns the expected prize amounts of $a$ and $\hat{a}$. More specifically, if $\hat{a}$ becomes more attractive relative to $b$ than $a$, it must have a weakly higher expected monetary payoff when evaluated at the beliefs of indifference between $a$ and $b$.
We begin by specifying that there are just two states of the world and imposing that Alice is risk averse. Two easy sufficient conditions are that $\hat{a}$ dominates $a$ or that $\hat{a}$ dominates $b$. The first cannot make $\hat{a}$ less attractive than $a$, no matter Alice’s belief; whereas the second means that Alice always prefers $\hat{a}$ to $b$. These conditions are quite strong, so we endeavor to find something weaker. We discover that when there is no such dominance improvement, the types of modification to $a$ that make $b$ less attractive depend crucially on the monotonicity of the decision problem and the monotonicity of the transformation.
When the decision problem is monotone in the sense that the state in which $b$ is preferred to $a$ (state $1$) yields higher payoffs to both $a$ and $b$, and the transformation is monotone in the sense that the payoff to $\hat{a}$ is also higher in state $1$, $\hat{a}$ becomes more attractive vis-a-vis $b$ than $a$ if $\hat{a}$’s payoffs are more spread out than $a$’s (or the dominance conditions hold). This is quite surprising, as this sort of spreading out of payoffs is typically disliked by risk-averse agents. What is happening is that increased risk aversion makes $a$ increasingly attractive in relation to $b$, viz., pushes the indifference belief to the right (toward state $1$). But this raises the expected monetary payoff to action $\hat{a}$, which drives the result.
If the decision problem is monotone but the transformation is not monotone, we obtain the stark finding that no non-dominance shift can make $\hat{a}$ better than $b$ for a larger set of beliefs than $a$. The logic is essentially the same as in the previous paragraph–risk aversion shifts the $a$-$b$ indifference belief to the right, but state $1$ is the worst state for action $\hat{a}$.
If the two-state decision problem is non-monotone, i.e., state 00 is better than state $1$ if and only if Alice chooses $a$, the results mirror the monotone environment. Now, if the transformation is monotone, $\hat{a}$ becomes more attractive vis-a-vis $b$ than $a$ if $\hat{a}$’s payoffs are more contracted than $a$’s (or the dominance conditions hold). On the other hand, any non-monotone transformation (provided the aforementioned expectation condition holds) makes $\hat{a}$ more attractive than $a$.
We then argue that understanding the two-state environment is key to understanding the general environment. Namely, in our first main theorem, Theorem 3.7, we argue that in the general (arbitrary number of states) environment, the necessary and sufficient conditions for the set of beliefs at which $\hat{a}$ is preferred to $b$ to be a superset of the set of beliefs at which $a$ is preferred to $b$ are simply the two-state conditions, aggregated. That is, it is necessary and sufficient that for any pair of states for which $a$ is strictly preferred in one and $b$ in the other, the two-state conditions hold for that pair of states.
If we do not assume Alice to be risk averse, our other main result (Theorem 3.9) is essentially a negative one: the set of beliefs for which $\hat{a}$ is preferred to $b$ is a superset of those for which $a$ is preferred to $b$, no matter Alice’s increasing utility function, if and only if $\hat{a}$ dominates $a$ or $b$. Concavity of $u$ already imposes significant structure on what transformations are improving, and convexity does too, but in the opposite manner. Combining these requirements leaves only dominance as necessary (and it is obviously sufficient).
1.1 Related Work
The body of work studying decision-making under uncertainty is sizeable. The work closest to this one is Pease and Whitmeyer (2023). There, we formulate a binary relation between actions: action $a$ is safer than $b$ if the the set of beliefs at which $a$ is preferred to $b$ grows larger, in a set inclusion sense, when we make Alice more risk averse. We use this characterization a number of times in the proofs in the current paper, and the extreme-points proof approach of our main theorems mirrors that that we take in proving our main theorem in the “Safety” paper.
Rothschild and Stiglitz (1970) is a seminal work that characterizes (mean-preserving) transformations of lotteries that are preferred by all risk-averse agents. Aumann and Serrano (2008) formulate an “measure of riskiness” of gambles, as do Foster and Hart (2009) (who are subsequently followed up upon by Bali, Cakici, and Chabi-Yo (2011) and Riedel and Hellmann (2015)). Crucially, these indices and measures correspond to inherently stochastic objects–the lotteries at hand. Our conception of an improvement to an action concerns comparisons of state-dependent payoffs, which are themselves non-random objects (they are just real numbers).
Naturally this work is also connected to the broader literature studying actions that are comparatively friendly toward risk. In addition to the aforementioned paper of ours, Pease and Whitmeyer (2023), which, like this one, centers around a decision-maker’s state-dependent payoffs to actions, this literature includes Hammond III (1974), Lambert and Hey (1979), Karlin and Novikoff (1963), Jewitt (1987), and Jewitt (1989). Notably, they are statements about lotteries, viz., random objects.
Ours is vaguely a comparative statics work–we’re changing an aspect of a decision problem and seeing how it affects a decision-maker’s choice. Our robustness criterion as well as the simplicity of our setting distinguishes our work from the standard pieces, e.g., Milgrom and Shannon (1994), Edlin and Shannon (1998), and Athey (2002). The works involving aggregation (Quah and Strulovici (2012), Choi and Smith (2017), and Kartik, Lee, and Rappoport (2023)) are closer still–as this inherently corresponds to distributional robustness–but none leave as free parameters both the distribution over states and the DM’s utility function, as we do. Special mention is due to Curello and Sinander (2019), who conduct a robust comparative statics exercise in which an analyst, with limited knowledge of an agent’s preferences, predicts the agent’s choice across menus.
2 Model
There is a topological space of states, $\Theta$, which is endowed with the Borel $\sigma$-algebra, and which we assume to be compact and metrizable. $\theta$ denotes a generic element of $\Theta$. We denote the set of all Borel probability measures on $\Theta$ by $\Delta\equiv\Delta\left(\Theta\right)$. There is also a decision-maker (DM), who is endowed with two actions. $A$ denotes the set of actions, and each action $a\in A$ is a continuous function from the state space to the set of outcomes, $a\colon\Theta\to\mathbb{R}$. Given a probability distribution over states $\mu\in\Delta$, an action is merely a (simple) lottery.
The DM is an expected-utility maximizer, with a von Neumann-Morgenstern utility function defined on the outcome space $u\colon\mathbb{R}\to\mathbb{R}$. We posit that $u$ is strictly increasing, weakly concave, and continuous. We further specify that no action $a\in A$ is weakly dominated by another. On occasion, we will drop the assumption that $u$ is weakly concave, merely requiring that it be strictly increasing (and continuous). For convenience, for any $a\in A$, we write $a_{\theta}\equiv a\left(\theta\right)$.
Given two actions $\left(a,b\right)$, we are interested in transformations of $a$ that affect the DM’s choice in a particular way. That is, we will modify $a$ to some new $\hat{a}$, in which case the new pair is $\left(\hat{a},b\right)$. To that end, we define the set $P_{b}\left(a\right)$ to be the subset of the probability simplex on which action $a$ is weakly preferred to $b$; formally,
$P_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\right\}\text{.}$
$\bar{P}_{b}\left(a\right)$, in turn, is the subset of the probability simplex on which $a$ is strictly preferred to $b$:
$\bar{P}_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\right\}\text{.}$
When $a$ is transformed to $\hat{a}$, we define the analogous sets $P_{b}\left(\hat{a}\right)$ and $\bar{P}_{b}\left(\hat{a}\right)$.
Definition 2.1.
Action $\hat{a}$ is $b$-Superior to action $a$ if, no matter the risk-averse agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{b}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{b}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $b$.
To put differently, action $a$ is $b$-superior to action $\hat{a}$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, concave, and continuous $u$.
Definition 2.2.
Action $\hat{a}$ is $b$-Better than action $a$ if, , no matter the (not necessarily risk-averse) agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{b}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{b}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $b$
To put differently, action $a$ is $b$-better than action $\hat{a}$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, but not necessarily concave, continuous $u$.
3 Robust Improvements
3.1 Two States
When there are two states, 00 and $1$, assuming without loss of generality that $a_{0}>b_{0}$ and $b_{1}>a_{1}$, and letting $\mu\in\left[0,1\right]$ denote the probability that the state is $1$, the DM prefers action $a$ if and only if
$\begin{split}\left(1-\mu\right)u\left(b_{0}\right)+\mu u\left(b_{1}\right)&\leq\left(1-\mu\right)u\left(a_{0}\right)+\mu u\left(a_{1}\right)\quad\Leftrightarrow\\ \mu&\leq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(a_{1}\right)}\equiv\bar{\mu}_{u}\text{.}\end{split}$
We can represent any transformation of $a$ to $\hat{a}$ as $\hat{a}_{0}=a_{0}+\varepsilon$ and $\hat{a}_{1}=a_{1}+\eta$, where $\varepsilon,\eta\in\mathbb{R}$. Accordingly, the DM prefers action $\hat{a}$ if and only if
$\mu\leq\frac{u\left(a_{0}+\varepsilon\right)-u\left(b_{0}\right)}{u\left(a_{0}+\varepsilon\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(a_{1}+\eta\right)}\equiv\hat{\mu}_{u}\text{.}$
When the DM is risk-neutral ($u$ is affine) we drop the subscript and write the indifference beliefs as $\bar{\mu}$ and $\hat{\mu}$.
In order for $\hat{a}$ to be $b$-superior to $a$, we need $\hat{\mu}_{u}\geq\bar{\mu}_{u}$ and $\hat{a}$ to be strictly preferred to $b$ in state 00 for any strictly increasing concave $u$. The easiest sufficient condition is $\hat{a}_{0}>b_{0}$ and $\hat{a}_{1}\geq a_{1}$, i.e., that $\hat{a}$ dominates $b$ with a strict preference in state 00. This is obviously quite strong–can we find anything weaker? Suppose $\hat{a}_{1}<b_{1}$. One variety of $u$ under consideration is those that are affine. In that case, some simple algebra reveals that $\hat{\mu}\geq\bar{\mu}$ if and only if
$\varepsilon\geq-\eta\underbrace{\frac{a_{0}-b_{0}}{b_{1}-a_{1}}}_{\eqqcolon\gamma}\text{.}$
This is equivalent to
$\left(b_{1}-a_{1}\right)\left(\hat{a}_{0}-b_{0}\right)\geq\left(a_{0}-b_{0}\right)\left(b_{1}-\hat{a}_{1}\right)\text{.}$
Of course, $\hat{a}_{0}>b_{0}$ if and only if
$\varepsilon>-\left(a_{0}-b_{0}\right)\text{.}$
Evidently, (when $\hat{a}_{1}<b_{1}$) both conditions are necessary for $\hat{\mu}_{u}\geq\bar{\mu}_{u}$ for all concave and strictly increasing $u$.
Scrutinizing Inequality 3.1, we see that it cannot be that both $\eta$ and $\varepsilon$ are negative (with at least one strict). Conversely, both Inequalities 3.1 and 3.1 are satisfied when both $\eta$ and $\varepsilon$ are positive. What about when the signs are opposed? When $\varepsilon=-\eta\gamma$, the expected values of actions $a$ and $\hat{a}$ are the same when evaluated at $\bar{\mu}$, i.e., $\bar{\mu}=\hat{\mu}$. Furthermore, when $\varepsilon=-\eta\gamma$ and $\hat{a}_{1},\hat{a}_{0}\in\operatorname*{conv}\left\{a_{0},a_{1}\right\}$–i.e., $\hat{a}_{1}$ and $\hat{a}_{0}$ lie in the convex hull of $a_{0}$ and $a_{1}$–the payoff lottery corresponding to action $\hat{a}$ and belief $\bar{\mu}$ is a mean-preserving contraction (MPC) of the lottery corresponding to $\bar{\mu}$ and $a$.
This line of reasoning suggests our first definition.
Definition 3.1.
Given $\left(a,b\right)$, action $\hat{a}$ is a $\bar{\mu}-$MPC of $a$ if $\varepsilon=-\eta\gamma$ and $\hat{a}_{1},\hat{a}_{0}\in\operatorname*{conv}\left\{a_{0},a_{1}\right\}$.
If $\hat{a}$ is a $\bar{\mu}-$MPC of $a$, then, in the terminology of Pease and Whitmeyer (2023), $\hat{a}$ is safer than $a$, i.e., becomes comparatively more attractive as the DM becomes more risk averse. We also define the opposite transformation; now, the payoff lottery corresponding to action $\hat{a}$ and belief $\bar{\mu}$ is a mean-preserving spread (MPS) of the lottery corresponding to $\bar{\mu}$ and $a$:
Definition 3.2.
Given $\left(a,b\right)$, action $\hat{a}$ is a $\bar{\mu}-$MPS of $a$ if $\varepsilon=-\eta\gamma$ and $a_{1},a_{0}\in\operatorname*{conv}\left\{\hat{a}_{0},\hat{a}_{1}\right\}$.
In this case, $a$ is safer than $\hat{a}$. One final useful definition is (statewise) dominance:
Definition 3.3.
Action $\hat{a}$ Dominates action $\tilde{a}$ if $\hat{a}_{\theta}\geq\tilde{a}_{\theta}$ for all $\theta\in\left\{0,1\right\}$.
Our first result involves the case in which the transformation from $a$ to $\hat{a}$ is Consistent: if $a_{0}>a_{1}$ then $\hat{a}_{0}\geq\hat{a}_{1}$, and if $a_{0}<a_{1}$ then $\hat{a}_{0}\leq\hat{a}_{1}$.
Lemma 3.4.
Suppose the transformation from $a$ to $\hat{a}$ is consistent. Then,
  1. 1.
    If $a_{0}>a_{1}$, $\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and $\hat{a}$ dominates $b$ or $\hat{a}$ dominates a $\bar{\mu}-$MPC of $a$.
  2. 2.
    If $a_{0}<a_{1}$, $\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and $\hat{a}$ dominates $b$ or $\hat{a}$ dominates a $\bar{\mu}-$MPS of $a$.
If either condition holds, we say that $\hat{a}$ is a Consistent improvement to $a$.
Proof.
The proof of this result, and all others omitted from the main text, may be found in the appendix (A).∎
The second part of this lemma is, on first glance, extremely surprising. How could it be that increasing the riskiness of an action makes the action more desirable to a risk-averse decision-maker? The first answer is that we are not quite increasing the riskiness of the action. As neither $a$ nor $b$ dominate each other, and as the reward from each is monotone in the state ($1$ is better than 00), action $a$ is safer (Pease and Whitmeyer (2023)) than $b$. That is, the set of beliefs at which the DM prefers $a$ to $b$ increases in the DM’s risk aversion. More specifically, the indifference belief, $\bar{\mu}_{u}$, must be to the right of the risk-neutral DM’s indifference belief $\bar{\mu}$.
When we make $\hat{a}$ more spread out than $a$, it is better in state $1$ than $a$. This interacts synergistically with an increased probability of state $1$. So, $a$ versus $\hat{a}$ is not the apples-to-apples comparison that it seems. In some sense, the more $u$ pushes the indifference belief between $a$ and $b$ to the right, the better it makes the $\bar{\mu}$-spread of $a$, $\hat{a}$ at this indifference belief. This effect dominates the negative effect of increased risk (with respect to $\bar{\mu}$), yielding the finding.
The first part is more mundane. $\hat{a}$ is less risky than $a$ when evaluated at the prior. Moreover, $\hat{a}$ is safer than $a$ so increased risk aversion makes $\hat{a}$ comparatively more attractive than $a$. In this light, it seems natural that $a$ being preferred to $b$ implies that $\hat{a}$ is preferred to $b$ too.
We now turn our attention to the scenario in which the transformation from $a$ to $\hat{a}$ is Inconsistent: if $a_{0}>a_{1}$ then $\hat{a}_{0}<\hat{a}_{1}$, and if $a_{0}<a_{1}$ then $\hat{a}_{0}>\hat{a}_{1}$.
Lemma 3.5.
Suppose the transformation from $a$ to $\hat{a}$ is inconsistent.
  1. 1.
    If $a_{0}>a_{1}$, then $\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and either
    1. (a)
      $\hat{a}$ dominates $b$, or
    2. (b)
      $\hat{a}$ does not dominate $b$ and Inequality 3.1 holds.
  2. 2.
    If $a_{0}<a_{1}$, $\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and $\hat{a}$ dominates $b$ or $a$.
If either condition holds, we say that $\hat{a}$ is a Inconsistent improvement to $a$.
This lemma is also unexpected. The first part says that if $a_{0}>a_{1}$ then as long as the expected payoff (in money) to $\hat{a}$ at the risk-neutral indifference belief $\bar{\mu}$ does not decrease, if $\hat{a}$ is a inconsistent transformation of $a$, it must be better versus $b$ than $a$. The second part is as negative as the first part is positive: if $a_{0}<a_{1}$, it is only if $\hat{a}$ dominates either $b$ or $a$ that the transformation makes $b$ less attractive.
Lemma 3.6.
Let $a_{0}=a_{1}$. Then, $\hat{a}$ is $b$-superior to $a$ if and only if either
  1. 1.
    $\hat{a}_{0}>b_{0}$ and $\hat{a}$ dominates $b$, or
  2. 2.
    $\hat{a}_{0}>b_{0}$, $b_{1}>\hat{a}_{1}\geq a_{1}$ and Inequality 3.1 holds.
If either condition holds, we say that $\hat{a}$ is a $b$-State improvement to $a$.
This lemma is more-or-less a modification to the second part of Lemma 3.4 and the logic is analogous–a particular variety of spread is necessary and sufficient for $\hat{a}$ to be $b$-superior to $a$.
We term $\hat{a}$ an Improvement to $a$ if $\hat{a}_{0}>b_{0}$ and $\hat{a}$ is either a consistent, an inconsistent, or a $b$-state improvement to $a$. Combining the three lemmas produces.
Theorem 3.7.
$\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}$ is an improvement to $a$.
Another perspective on this theorem is a purely geometric one. Take the line, in belief $\mu$,
$\ell_{a}(\mu)=a_{0}(1-\mu)+a_{1}\mu\text{,}$
which is the expected payoff (in money) of action $a$ as a function of the belief $\mu\in\left[0,1\right]$. In turn, lines
$\ell_{\hat{a}}(\mu)=\hat{a}_{0}(1-\mu)+\hat{a}_{1}\mu\quad\text{and}\quad\ell_{b}(\mu)=b_{0}(1-\mu)+b\mu\text{,}$
are the expected payoffs to actions $\hat{a}$ and $b$, respectively. It turns out that $b$-superiority of $\hat{a}$ is equivalent to $\ell_{\hat{a}}$ lying weakly above a counter-clockwise rotation of $\ell(a)$ around the point $\left(\bar{\mu},\ell_{a}\left(\bar{\mu}\right)\right)$, provided the rotation isn’t too extreme so that $\hat{a}_{0}\leq b_{0}$. When $\ell_{\hat{a}}$ is a counter-clockwise rotation of $\ell_{a}$, we have made it “more like $\ell_{b}$,” and this cannot hurt $\hat{a}$’s comparative attractiveness versus $b$.
Figure 1 illustrates the benefits of a counter-clockwise rotation. The top three lines are the expected payoffs, in belief $\mu$, to $b$, $\hat{a}$, and $a$; in dashed red, dotted green, and solid blue, respectively. The bottom three lines are the expected utilities (with $u\left(\cdot\right)=\sqrt{1+\cdot}-3$) to the three actions, maintaining the same color and fill correspondences. Line $x=\bar{\mu}_{u}$ is in (thin) solid black and $x=\hat{\mu}_{u}$ is in (thin) dotted orange. Here is an interactive link, where the line can be rotated by moving the $l$ slider. See how $\hat{\mu}_{u}$ moves!
Figure 1: Improving $a$. Try it yourself (rotate by moving the slider)!
Before going on, let us jot down this informal discussion as a formal result.
Corollary 3.8.
$\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and there exists a $\lambda\in\left[0,1\right]$ such that $\ell_{\hat{a}}$ lies weakly above the line $\lambda\ell_{a}+\left(1-\lambda\right)\ell_{b}$.
So far we have assumed mild but meaningful structure on the DM’s utility function. More is better ($u$ is strictly increasing in the reward), and the DM is risk-averse ($u$ is, at least weakly, concave). Suppose we remove the assumption of risk aversion. We still assume that $u$ is strictly increasing but now allow it to take arbitrary shape.
As we now document, this is, in a sense, a step too far. Without the assumption of risk aversion, the only transformations that must make $b$ less attractive are unambiguous dominance improvements from $a$ to $\hat{a}$. This is intuitive: think back to Theorem 3.7. There, the necessary transformations were either dominance shifts (which obviously lead $\hat{a}$ to be better versus $b$ than $a$) or spreads or contractions of $a$, depending on the initial structure of the problem. However, the opposite modifications are needed when $u$ is convex–which is now permitted. Consequently, other than a dominance improvement, there is no way to ensure the desired outcome when the DM may be either risk loving or risk averse. Formally,
Theorem 3.9.
Action $\hat{a}$ is $b$-better than action $a$ if and only if $\hat{a}_{0}>b_{0}$ and $\hat{a}$ dominates $a$ or $b$.
3.2 More Than Two States
Now let $\Theta$ be an compact and metrizable space (endowed with the Borel $\sigma$-algebra). We partition $\Theta$ into three sets:
$\begin{split}\mathcal{A}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>b_{\theta}\right\}\quad\mathcal{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=b_{\theta}\right\}\text{.}\end{split}$
We extend the earlier (two-state) definition of dominance:
Definition 3.10.
Action $\hat{a}$ Dominates action $\tilde{a}$ if $\hat{a}_{\theta}\geq\tilde{a}_{\theta}$ for all $\theta\in\Theta$.
Then, for any pair $\theta\in\mathcal{A}$ and $\theta^{\prime}\in\mathcal{B}$, let $\bar{\mu}^{\theta,\theta^{\prime}}_{u}$ denote the binary distribution over states supported on only $\theta$ and $\theta^{\prime}$ at which the DM is indifferent between $a$ and $b$, i.e.,
$\bar{\mu}_{u}^{\theta,\theta^{\prime}}\coloneqq\frac{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)}{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)+u\left(b_{\theta^{\prime}}\right)-u\left(a_{\theta^{\prime}}\right)}\text{.}$
For any pair of states $\theta\in\mathcal{A}$ and $\theta^{\prime}\in\mathcal{B}$, we define an X improvement–where X is either “consistent,” “inconsistent,” or “$b$-state”–to be exactly as defined in the two-state setting (§$\S$3.1), where state 00 is replaced with $\theta$, $1$ with $\theta^{\prime}$, and $\bar{\mu}_{u}$ with $\bar{\mu}_{u}^{\theta,\theta^{\prime}}$. Likewise, we extend the definition of “improvement,” so that $\hat{a}$ is an improvement to $a$ if for any $\theta^{\dagger}\in\mathcal{C}$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$, and for any $\theta\in\mathcal{A}$ and $\theta^{\prime}\in\mathcal{B}$, $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ is either a consistent, an inconsistent, or a $b$-state improvement to $a$. Then,
Theorem 3.11.
$\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}$ is an improvement to $a$.
Proof.
We merely sketch the proof, as this theorem is proved identically to Theorem 3.4 in Pease and Whitmeyer (2023). The key observation is that $P_{b}(a)$ ($P_{b}(\hat{a})$) is a compact convex subset of $\Delta$ whose extreme points are the degenerate probability distributions at which $a$ ($\hat{a}$) is preferred to $b$ and the binary distributions at which the DM is indifferent between $a$ ($\hat{a}$) and $b$. By the Krein-Milman theorem, these sets of extreme points characterize the sets $P_{b}(a)$ and $P_{b}(\hat{a})$, so the result is equivalent to the extreme points acting in a particular way (moving “away” from each $\theta\in\mathcal{A}$ along each edge of the simplex connecting $\theta\in\mathcal{A}$ and $\theta^{\prime}\in\mathcal{B}$) as we go from $a$ to $\hat{a}$. Hence, it is necessary and sufficient to simply aggregate the two-state result. ∎
This extreme points argument goes through without alteration to produce our second result.
Theorem 3.12.
Action $\hat{a}$ is $b$-better than action $a$ if and only if for all $\theta\in\mathcal{A}$ $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ dominates $a$ or $b$.
Proof.
Omitted as the already-cited extreme-points argument is valid. ∎
4 One Versus Many
What if we look for improvements to $a$ that make it more attractive in comparison to multiple alternatives, not just a single one? For simplicity, we restrict our attention to the two-state environment. Let $A$ be the DM’s finite set of actions, where we assume that none are (strictly) dominated. Indexing the actions by a finite set $I\subset\mathbb{N}$ (with $\left|I\right|=m\geq 3$), we assume without loss of generality (as we could just relabel actions) that for all $i\in I$, $a_{0}^{i}>a_{0}^{j}$ and $a_{1}^{i}<a_{1}^{j}$ for all $j>i$ ($j\in I$).
There are two special actions, $a^{1}$ and $a^{m}$. The first is uniquely optimal in state 00 and the second is uniquely optimal in state $1$. We term these the two extreme actions. For a specified $a\in A$, we define $B\coloneqq A\setminus\left\{a\right\}$ and
$P_{B}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\max_{a^{j}\in B}\mathbb{E}_{\mu}u\left(a^{j}_{\theta}\right)\right\}\text{.}$
We define $\bar{P}_{B}\left(a\right)$ to be the analogous set with a strict inequality, and sets $P_{B}\left(\hat{a}\right)$ and $\bar{P}_{B}\left(\hat{a}\right)$ their compatriots once we’ve replaced $a$ with $\hat{a}$.
Definition 4.1.
Action $\hat{a}$ is $B$-Superior to action $a$ if the set of beliefs at which the DM prefers $a$ to the actions other than $a$, $P_{B}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to those other actions, $P_{B}(\hat{a})$; and $\bar{P}_{B}(a)\subseteq\bar{P}_{B}(\hat{a})$; no matter the risk-averse agent’s utility function.
We have the following stark result.
Proposition 4.2.
If $a=a^{1}$ ($a=a^{m}$), $\hat{a}$ is $B$-superior to $a$ if and only if $a$ is $b$-superior to $a^{2}$ ($a^{m-1}$). If $a$ is not one of the extreme actions, $\hat{a}$ is $B$-superior to $a$ if and only if $\hat{a}$ dominates $a$.
Proof.
[…]
If $a\neq a^{1},a^{m}$ it is immediate that $\hat{a}$ dominating $a$ implies $P_{B}(a)\subseteq P_{B}(\hat{a})$.
[…]
The logic behind this proposition is straightforward. For an action to improve versus the action to its right, we need the induced (risk-neutral) line in belief space to lie above some counterclockwise rotation around the indifference point between the two actions. For an action to improve versus the action to its left, we need the line in belief space to lie above the indifference point between the two actions (which is strictly to the left of the other indifference point). A dominance shift is the only way to reconcile these requirements.
5 Robustness to Aggregate Risk
What if the DM also has unknown (to us) random wealth? Formally, in any state $\theta$, the DM’s utility from action $a$ is $\mathbb{E}_{H}u\left(a_{\theta}+W\right)$, where $W$ is a state-independent (finite-mean) random variable .
Abusing notation, we define $P_{a,b}(a)$ to be the subset of $\Delta$ for which $a$ is preferred to $b$:
$P_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu,H}u\left(a_{\theta}+W\right)\geq\mathbb{E}_{\mu,H}u\left(b_{\theta}+W\right)\right\}\text{.}$
$\bar{P}_{b}\left(a\right)$ is this set’s strict-preference counterpart.
We extend our definitions in the natural way:
Definition 5.1.
Action $\hat{a}$ is $b$- and Wealth-Superior to action $a$ if the set of beliefs at which the DM prefers $a$ to $b$, $P_{b}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to $b$, $P_{b}(\hat{a})$; and the set of beliefs at which the DM strictly prefers $a$ to $b$, $\bar{P}_{b}(a)$, is a subset of the set of beliefs at which the DM strictly prefers $\hat{a}$ to $b$, $\bar{P}_{b}(\hat{a})$; no matter the risk-averse agent’s utility function (for any strictly increasing, concave $u$) and no matter the agent’s state-independent random wealth $W$ (for any finite-mean, real-valued random variable $W$).
Definition 5.2.
Action $\hat{a}$ is $b$- and Wealth-Better than action $a$ if the set of beliefs at which the DM prefers $a$ to $b$, $P_{b}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to $b$, $P_{b}(\hat{a})$; and the set of beliefs at which the DM strictly prefers $a$ to $b$, $\bar{P}_{b}(a)$, is a subset of the set of beliefs at which the DM strictly prefers $\hat{a}$ to $b$, $\bar{P}_{b}(\hat{a})$; no matter the (not necessarily risk-averse) agent’s utility function (for any strictly increasing, concave $u$) and no matter the agent’s state-independent random wealth $W$ (for any finite-mean, real-valued random variable $W$).
Then,
Theorem 5.3.
$\hat{a}$ is $b$- and wealth-superior to $a$ if and only if $\hat{a}$ is an improvement to $b$. $\hat{a}$ is $b$- and wealth-better than $a$ if and only if for all $\theta\in\mathcal{A}$ $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ dominates $a$ or $b$.
6 Everything Everywhere All at Once
In this paper, we focus on altering a single action, leaving the other action (or actions) unaltered. This is reasonable in many situations–for example, in the two applications in the next section, the other action is the DM’s outside option, which is unalterable. But what if we change both actions $a$ and $b$? As we will see, the basic intuition from the earlier results remains true, though we only provide a partial analysis, primarily to get bogged down in a tedium of cases. We also restrict attention to the two-state environment, as we could just follow the earlier approach to generate the result for a general state space (it will again just be the aggregation of the two-state conditions).
Now, we are going from $a$ to $\hat{a}$ and $b$ to some $\hat{b}$. To that end, we define
$P_{\hat{b}}\left(\hat{a}\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(\hat{b}_{\theta}\right)\right\}\text{,}$
understanding $\bar{P}_{\hat{b}}\left(\hat{a}\right)$ to be this set’s strict inequality counterpart.
Definition 6.1.
Given $a$ and $b$, action $\hat{a}$ is $\hat{b}$-Superior to action $a$ if, no matter the risk-averse agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{\hat{b}}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{\hat{b}}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $\hat{b}$.
It is easy to obtain an analogous inequality to Inequality 3.1, which is necessary for $\hat{b}$-superiority when $\hat{a}_{1}<\hat{b}_{1}$:
$\left(b_{1}-a_{1}\right)\left(\hat{a}_{0}-\hat{b}_{0}\right)\geq\left(a_{0}-b_{0}\right)\left(\hat{b}_{1}-\hat{a}_{1}\right)\text{.}$
Obviously, $\hat{a}$ dominating $a$ or $\hat{b}$ and $\hat{b}$ being dominated by $b$ or $\hat{a}$ imply $\hat{b}$-superiority. Here is a less obvious sufficient condition.
Proposition 6.2.
If Inequality 6 holds, $a_{1}\leq\hat{a}_{1}<b_{1}\leq\hat{b}_{1}$, and $\hat{b}_{0}\leq b_{0}<\hat{a}_{0}\leq a_{0}$, $\hat{a}$ is $\hat{b}$-superior to action $a$.
As we discussed above, $\hat{a}$ improving upon $a$ versus $b$ (when $b$ is fixed) is equivalent to $\hat{a}$’s expected payoff in belief space lying above a counter-clockwise rotation of $a$’s payoff. When we alter both actions, the rotation point may now shift–think vaguely of a co-movement of payoffs to both $a$ and $b$–and $\hat{a}$’s payoff and $\hat{b}$’s payoffs are both counter-clockwise rotations about this point.
7 Applications
7.1 Robustly Optimal Bilateral Trade Modifications
There is a buyer ($B$) and a seller ($S$). $B$ possesses an asset that pays out $H>0$ in state $1$ and 00 in state 00. The state is contractible and the status quo trade agreement sees transfers of size $\gamma_{\theta}\in\mathbb{R}$ from $B$ to $S$ in each state $\theta\in\Theta$. We assume that $\gamma_{0}>0$ and $0<\gamma_{1}<H$.
$B$’s payoff from transacting is $-\gamma_{0}$ in state 00 and $H-\gamma_{1}$ in state $1$. Similarly, $S$’s payoff from transacting is $\gamma_{0}$ in state 00 and $\gamma_{1}$ in state $1$. $B$’s outside option is the sure-thing 00 and $S$’s outside option is the asset, with random payoff $H$ in state $1$ and 00 in state 00.
We suppose that the status-quo is acceptable–given the agents’ subjective beliefs about uncertainty and their (concave) utility functions, each is willing to participate in the arrangement. What modifications to the arrangement are robustly optimal in that $B$ and $S$ will still remain willing to participate?
Observe that if $B$ and $S$ are risk-neutral, their indifference belief is
$\bar{\mu}=\frac{\gamma_{0}}{H-\gamma_{1}+\gamma_{0}}\text{,}$
and $S$ ($B$) trades whenever she is sufficiently pessimistic (optimistic). Furthermore, in any modification, it is clear that Inequality 3.1 must hold with equality for both $B$ and $S$, as transfers must balance. There are two cases to consider: 1. $\gamma_{0}<\gamma_{1}$, and 2. $\gamma_{0}>\gamma_{1}$.
Case 1. Suppose $\gamma_{0}<\gamma_{1}$. By Lemma 3.4, a monotone alteration to the trade agreement–one that is such that $\hat{\gamma}_{0}\leq\hat{\gamma_{1}}$–is acceptable to $S$ if and only it is a spread of the initial agreement:
$0<\hat{\gamma_{0}}\leq\gamma_{0}\quad\text{and}\quad\hat{\gamma}_{1}=\frac{\left(\gamma_{0}-\hat{\gamma}_{0}\right)\left(1-\bar{\mu}\right)+\gamma_{1}\bar{\mu}}{\bar{\mu}}\text{.}$
Similarly, a monotone alteration–one that is such that $-\hat{\gamma}_{0}\leq H-\hat{\gamma}_{1}$–is acceptable to $B$ if and only if it is a contraction of the initial agreement, which corresponds to the conditions in Expression 7.1. By Lemma 3.5, there is no (robustly) acceptable non-monotone alteration.
Remark 7.1.
If $\gamma_{0}<\gamma_{1}$, the set of acceptable alterations are monotone and satisfy the conditions in Expression 7.1.
Case 2. Now suppose $\gamma_{0}>\gamma_{1}$. Now, the only monotone alteration to the trade agreement acceptable to $S$ is a contraction of the initial agreement:
$\frac{\gamma_{0}H}{H+\gamma_{0}-\gamma_{1}}\leq\hat{\gamma_{0}}\leq\gamma_{0}\quad\text{and}\quad\hat{\gamma}_{1}=\frac{\left(\gamma_{0}-\hat{\gamma}_{0}\right)\left(1-\bar{\mu}\right)+\gamma_{1}\bar{\mu}}{\bar{\mu}}\text{.}$
As before, the only monotone alteration $B$ is guaranteed to accept is a contraction, so Expression 7.1 specifies the binding constraints.
Finally, Lemma 3.5 implies that $S$ and $B$ will accept any non-monotone alteration (that preserves the expected value) and strict optimality in state 00 and $1$, respectively:
$0<\hat{\gamma_{0}}<\frac{\gamma_{0}H}{H+\gamma_{0}-\gamma_{1}}\quad\text{and}\quad\hat{\gamma}_{1}=\frac{\left(\gamma_{0}-\hat{\gamma}_{0}\right)\left(1-\bar{\mu}\right)+\gamma_{1}\bar{\mu}}{\bar{\mu}}\text{.}$
Remark 7.2.
If $\gamma_{0}<\gamma_{1}$, the set of acceptable alterations either satisfy the conditions in Expression 7.1 (monotone) or Expression 7.1 (non-monotone).
7.2 Robustly Optimal Insurance Modifications
Now consider the scenario of a risk-neutral insurer and a consumer. The consumer’s payoff without insurance is $H>0$ in state $1$ and 00 in state 00. The status-quo insurance policy pays out $\alpha_{0}>0$ to the consumer in state 00 and $-\alpha_{1}$ ($H>\alpha_{1}>0$) in state $1$. We assume further that $\alpha_{0}>H-\alpha_{1}$. Thus, if the consumer buys the status-quo contract, she gets $\alpha_{0}>0$ in state 00 and $H-\alpha_{1}>0$ in state $1$. In contrast, if she does not buy insurance, she gets 00 and $H$ in the two states, respectively.
The consumer’s (risk-neutral) indifference belief is
$\bar{\mu}=\frac{\alpha_{0}}{\alpha_{0}+\alpha_{1}}\text{.}$
A new contract is acceptable if the consumer is willing to accept it, given that she is willing to accept the status-quo contract, no matter her concave utility function or prior. Appealing to Lemma 3.5, we have
Remark 7.3.
If the new contract $\left(\hat{\alpha}_{0},\hat{\alpha}_{1}\right)$ is such that $\hat{\alpha}_{0}/\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}\right)\geq\alpha_{0}/\left(\alpha_{0}+\alpha_{1}\right)$, $\hat{a}_{0}>\alpha_{0}$, and $\hat{\alpha}_{1}\leq\hat{\alpha}_{0}\alpha_{1}$, the new contract is acceptable.
Appendix A Omitted Proofs
A.1 Lemma 3.4 Proof
Proof.
[…]
There are two cases: either 1. $u$ is such that $\bar{\mu}_{u}\geq\bar{\mu}$, or 2. $u$ is such that $\bar{\mu}_{u}<\bar{\mu}$.
Case 1. As the line (in belief space) corresponding to $\hat{a}$ crosses the line corresponding to $a$ from above and $\bar{\mu}_{u}\geq\bar{\mu}$, it must be that
$\bar{\mu}_{u}\hat{a}_{1}+\left(1-\bar{\mu}_{u}\right)\hat{a}_{0}\geq\bar{\mu}_{u}a_{1}+\left(1-\bar{\mu}_{u}\right)a_{0}\text{.}$
Moreover, in the terminology of Pease and Whitmeyer (2023), $\hat{a}$ is safer than $a$, so
$\begin{split}\bar{\mu}_{u}u\left(\hat{a}_{1}\right)+\left(1-\bar{\mu}_{u}\right)u\left(\hat{a}_{0}\right)&\geq\bar{\mu}_{u}u\left(a_{1}\right)+\left(1-\bar{\mu}_{u}\right)u\left(a_{0}\right)\\ &=\bar{\mu}_{u}u\left(b_{1}\right)+\left(1-\bar{\mu}_{u}\right)u\left(b_{0}\right)\text{,}\end{split}$
where the last line follows from the definition of $\bar{\mu}_{u}$. Consequently, the DM prefers $\hat{a}$ to $b$ for all $\mu\in\left[0,\bar{\mu}_{u}\right]$, strictly so for all $\mu\in\left[0,\bar{\mu}_{u}\right)$.
Case 2. Suppose now that $\bar{\mu}_{u}<\bar{\mu}$, and suppose for the sake of contradiction that
$\bar{\mu}_{u}u\left(\hat{a}_{1}\right)+\left(1-\bar{\mu}_{u}\right)u\left(\hat{a}_{0}\right)<\bar{\mu}_{u}u\left(b_{1}\right)+\left(1-\bar{\mu}_{u}\right)u\left(b_{0}\right)\text{.}$
But, as $a_{0}>\hat{a}_{0}>b_{0}$ and $b_{1}>\hat{a}_{1}>a_{1}$, this implies that $\hat{a}$ is strictly dominated when the DM has utility $u$, contradicting the two Proposition 1s in Weinstein (2016) and Battigalli, Cerreia-Vioglio, Maccheroni, and Marinacci (2016), which state that an increase in risk aversion cannot shrink the set of undominated actions.
Now let $a_{0}<a_{1}$. As before, it suffices to show the result when $\hat{a}$ is a $\bar{\mu}-$MPS of $a$, as subsequent dominance can only improve $\hat{a}$’s attractiveness in relation to $b$. This implies
$b_{0}<\hat{a}_{0}<a_{0}<a_{1}<\hat{a}_{1}<b_{1}\text{.}$
Consequently,
$\begin{split}\hat{\mu}_{u}=\frac{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)}{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(\hat{a}_{1}\right)}&=\frac{\frac{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)}{\hat{a}_{0}-b_{0}}}{\frac{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)}{\hat{a}_{0}-b_{0}}+\frac{u\left(b_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-b_{0}}}\\ &\geq\frac{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}}{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}+\frac{u\left(b_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-b_{0}}}\\ &=\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(a_{0}-b_{0}\right)\frac{u\left(b_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-b_{0}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(b_{1}-a_{1}\right)\frac{u\left(b_{1}\right)-u\left(\hat{a}_{1}\right)}{b_{1}-\hat{a}_{1}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(a_{1}\right)}=\bar{\mu}_{u}\text{,}\end{split}$
where the first inequality follows from the Three-chord lemma (Theorem 1.16 in Phelps (2009)), the second inequality from Inequality 3.1, and the third inequality from the Three-chord lemma.
[…]
We want to find a concave, strictly increasing $u(\cdot)$ such that $\hat{\mu}_{u}<\bar{\mu}_{u}$. Here is one:
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{0},\\ \iota x+(1-\iota)a_{0},\quad&\text{if}\quad x>a_{0}\text{,}\end{cases}$
for some $\iota\in(0,1)$. Directly, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+u(b_{1})-a_{1}}-\frac{a_{0}+\iota\varepsilon-b_{0}}{a_{0}+\iota\varepsilon-b_{0}+u(b_{1})-a_{1}-\eta}\text{.}$
This expression is continuous and decreasing in $\iota$ and is strictly positive when $\iota=0$. Accordingly for all sufficiently small $\iota>0$, we have that $\hat{\mu}_{u}<\bar{\mu}_{u}$.
Now let $a_{1}>a_{0}$ and suppose for the sake of contraposition that $\hat{a}$ neither dominates $b$ nor a $\bar{\mu}-$MPS of $a$. Again, this implies $b_{1}>\hat{a}_{1}$ and that $\hat{a}_{1}<a_{1}$ or $\hat{a}_{0}<a_{0}$. As we may assume Inequality 3.1 holds, it cannot be that both hold, which means that $\hat{a}_{1}<a_{1}$ and $\hat{a}_{0}>a_{0}$. Consequently, by monotonicity of the transformation, we must have
$b_{0}<a_{0}<\hat{a}_{0}\leq\hat{a}_{1}<a_{1}<b_{1}\text{.}$
Let
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{1},\\ \iota x+(1-\iota)a_{1},\quad&\text{if}\quad x>a_{1}\text{,}\end{cases}$
for some $\iota\in(0,1)$. Directly, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+\iota(b_{1}-a_{1})}-\frac{\hat{a}_{0}-b_{0}}{\hat{a}_{0}-b_{0}+\iota(b_{1}-a_{1})+a_{1}-\hat{a}_{1}}\text{,}$
which is strictly positive when $\iota=0$.∎
A.2 Lemma 3.5 Proof
Proof.
[…]
Now suppose $a_{0}>\hat{a}_{0}>b_{0}$, $b_{1}>\hat{a}_{1}$, and Inequality 3.1 holds. Fix $u$. By the non-monotonicity of the transformation
$b_{0}<\hat{a}_{0}<\hat{a}_{1}<b_{1}\text{,}$
viz., $\hat{a}$ is safer than $b$. Furthermore, this means that there exists a $\mu^{\prime}$ such that the DM prefers $b$ to $\hat{a}$ for all $\mu\geq\mu^{\prime}\in\left(0,1\right)$, strictly so for $\mu>\mu^{\prime}$. As Inequality 3.1 holds and $a_{0}>\hat{a}_{0}$, $\hat{a}_{1}>a_{1}$. This implies that there exists $\mu^{\prime\prime}\in\left(0,1\right)$ such that the DM prefers $a$ to $\hat{a}$ for all $\mu\leq\mu^{\prime\prime}$ strictly so for $\mu<\mu^{\prime\prime}$.
As $\hat{a}$ cannot be strictly dominated (appealing to the Proposition 1s in Weinstein (2016) and Battigalli et al. (2016)), $\mu^{\prime\prime}\leq\mu^{\prime}$. Consequently, $\hat{\mu}_{u}\geq\bar{\mu}_{u}$.
If $a_{0}<a_{1}$, this direction is immediate, as dominance over $a$ or $b$ makes $\hat{a}$ relatively more attractive versus $b$ than $a$ is.
[…]
So, let $a_{0}<a_{1}$. We assume $\hat{a}_{0}>b_{0}$ or else the result is immediate. Suppose for the sake of contradiction that $\hat{a}$ neither dominates $b$ nor $a$. This implies that $b_{1}>\hat{a}_{1}$. As $\hat{a}$ does not dominate $a$ it must be that $a_{1}>\hat{a}_{1}$ or else we would have $\hat{a}_{0}>\hat{a}_{1}\geq a_{1}>a_{0}$. We assume that Inequality 3.1 holds, or else we are done. This means that $\hat{a}_{0}>a_{0}$. Summarizing things, we have
$\hat{a}_{0}>a_{0}>b_{0},\ b_{1}>a_{1}>\hat{a}_{1},\ a_{0}<a_{1},\ b_{0}<b_{1},\text{ and }\hat{a}_{1}<\hat{a}_{0}\text{.}$
The location of $\hat{a}_{0}$ is ambiguous with respect to $a_{1}$, but that won’t matter.
Let
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{1},\\ \iota x+(1-\iota)a_{1},\quad&\text{if}\quad x>a_{1}\text{,}\end{cases}$
for some $\iota\in\left(0,1\right)$ in which case $\bar{\mu}-\hat{\mu}$ equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+\iota\left(b_{1}-a_{1}\right)}-\frac{u(\hat{a}_{0})-b_{0}}{u(\hat{a}_{0})-b_{0}+\iota(b_{1}-a_{1})+a_{1}-\hat{a}_{1}}\text{,}$
which is strictly positive when $\iota=0$.∎
A.3 Lemma 3.6 Proof
Proof.
[…]
$b_{1}>\hat{a}_{1}>a_{1}=a_{0}>\hat{a}_{0}>b_{0}\text{.}$
Then, the chain of inequalities in Expression A.1 is valid, yielding the result.
[…]
Take
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{0},\\ \iota x+(1-\iota)a_{0},\quad&\text{if}\quad x>a_{0}\text{,}\end{cases}$
for some $\iota\in(0,1)$, so that $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+\iota\left(b_{1}-a_{1}\right)}-\frac{a_{0}+\iota\varepsilon-b_{0}}{a_{0}+\iota\varepsilon-b_{0}+\iota\left(b_{1}-a_{1}\right)+a_{1}-\hat{a}_{1}}\text{,}$
which is strictly positive when $\iota=0$.∎
A.4 Theorem 3.9 Proof
Proof.
[…]
Case 1. First suppose $a_{1}\geq a_{0}$ so that we have the chain
$b_{0}<\hat{a}_{0}<a_{0}\leq a_{1}<\hat{a}_{1}<b_{1}\text{.}$
Take
$u(x)=\begin{cases}\iota x+(1-\iota)\hat{a}_{0},\quad&\text{if}\quad x\leq\hat{a}_{0},\\ x,\quad&\text{if}\quad x>\hat{a}_{0}\text{,}\end{cases}$
for some $\iota\in(0,1)$, so that $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-\hat{a}_{0}-\iota\left(b_{0}-\hat{a}_{0}\right)}{a_{0}-\hat{a}_{0}-\iota\left(b_{0}-\hat{a}_{0}\right)+b_{1}-a_{1}}-\frac{\iota\left(\hat{a}_{0}-b_{0}\right)}{\iota\left(\hat{a}_{0}-b_{0}\right)+b_{1}-\hat{a}_{1}}\text{,}$
which is strictly positive when $\iota=0$.
Now suppose $a_{1}<a_{0}$, in which case we have
$b_{0}<\hat{a}_{0}<a_{0},\quad a_{1}<\hat{a}_{1}<b_{1},\quad\text{and}\quad a_{1}<a_{0}\text{.}$
The same $u$ defined in Expression A.4 will do. Thus, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-\hat{a}_{0}-\iota\left(b_{0}-\hat{a}_{0}\right)}{a_{0}-\hat{a}_{0}-\iota\left(b_{0}-\hat{a}_{0}\right)+u(b_{1})-u(a_{1})}-\frac{\iota\left(\hat{a}_{0}-b_{0}\right)}{\iota\left(\hat{a}_{0}-b_{0}\right)+u(b_{1})-u(\hat{a}_{1})}\text{.}$
If $b_{1}\leq\hat{a}_{0}$, this equals
$1-\frac{\hat{a}_{0}-b_{0}}{\hat{a}_{0}-b_{0}+b_{1}-\hat{a}_{1}}>0\text{,}$
when $\iota=0$. If $b_{1}>\hat{a}_{0}$, it equals
$\frac{a_{0}-\hat{a}_{0}}{a_{0}-\hat{a}_{0}+b_{1}-u(a_{1})}-0>0\text{,}$
when $\iota=0$.
Case 2. First suppose $\hat{a}_{1}\geq\hat{a}_{0}$ so that we have the chain
$b_{0}<a_{0}<\hat{a}_{0}\leq\hat{a}_{1}<a_{1}<b_{1}\text{.}$
Take
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{1},\\ \iota x+(1-\iota)a_{1},\quad&\text{if}\quad x>a_{1}\text{,}\end{cases}$
for some $\iota\in(0,1)$, so that $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+\iota\left(b_{1}-a_{1}\right)}-\frac{\hat{a}_{0}-b_{0}}{\hat{a}_{0}-b_{0}+\iota\left(b_{1}-a_{1}\right)+a_{1}-\hat{a}_{1}}\text{,}$
which is strictly positive when $\iota=0$.
Now suppose $\hat{a}_{1}<\hat{a}_{0}$, in which case we have
$b_{0}<a_{0}<\hat{a}_{0},\quad\hat{a}_{1}<a_{1}<b_{1},\quad\text{and}\quad\hat{a}_{1}<\hat{a}_{0}\text{.}$
Using the $u$ from Expression A.4, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\iota\left(b_{1}-a_{1}\right)}-\frac{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)}{u\left(\hat{a}_{0}\right)-u\left(b_{0}\right)+\iota\left(b_{1}-a_{1}\right)+a_{1}-\hat{a}_{1}}\text{.}$
If $b_{0}\geq a_{1}$, this equals
$\frac{a_{0}-b_{0}}{a_{0}-b_{0}+b_{1}-a_{1}}-0>0\text{,}$
when $\iota=0$. If $b_{0}<a_{1}$, this equals
$1-\frac{u\left(\hat{a}_{0}\right)-b_{0}}{u\left(\hat{a}_{0}\right)-b_{0}+a_{1}-\hat{a}_{1}}>0\text{,}$
when $\iota=0$. ∎
A.5 Theorem 5.3 Proof
Proof.
[…]
We prove a sequence of lemmas. We note that $\bar{\mu}$ and $\hat{\mu}$ remain unchanged (the aggregate risk cancels out), but redefine
$\bar{\mu}_{u}\coloneqq\frac{v\left(a_{0}\right)-\mathbb{E}_{H}v\left(b_{0}\right)}{v\left(a_{0}\right)-\mathbb{E}_{H}u\left(b_{0}+W\right)+v\left(b_{1}\right)-\mathbb{E}_{H}u\left(a_{1}+W\right)}\text{,}$
and
$\hat{\mu}_{u}\coloneqq\frac{v\left(\hat{a}_{0}\right)-\mathbb{E}_{H}u\left(b_{0}+W\right)}{v\left(\hat{a}_{0}\right)-\mathbb{E}_{H}u\left(b_{0}+W\right)+v\left(b_{1}\right)-\mathbb{E}_{H}u\left(\hat{a}_{1}+W\right)}\text{.}$
Lemma A.1.
Suppose the transformation from $a$ to $\hat{a}$ is monotone. Then,
  1. 1.
    If $a_{0}>a_{1}$, $\hat{a}$ is $b$- and wealth-superior to $a$ if $\hat{a}$ dominates a $\bar{\mu}-$MPC of $a$.
  2. 2.
    If $a_{0}<a_{1}$, $\hat{a}$ is $b$- and wealth-superior to $a$ if $\hat{a}$ dominates a $\bar{\mu}-$MPS of $a$.
Proof.
We follow the proof of Lemma 3.4. There are two cases: either 1. $u$ is such that $\bar{\mu}_{u}\geq\bar{\mu}$, or 2. $u$ is such that $\bar{\mu}_{u}<\bar{\mu}$.
Case 1. As $\hat{a}_{1}>a_{1}$ and $\bar{\mu}_{u}\geq\bar{\mu}$, it must be that
$\bar{\mu}_{u}\hat{a}_{1}+\left(1-\bar{\mu}_{u}\right)\hat{a}_{0}\geq\bar{\mu}_{u}a_{1}+\left(1-\bar{\mu}_{u}\right)a_{0}\text{.}$
Moreover, in the terminology of Pease and Whitmeyer (2023), $\hat{a}$ hedges risk better than $a$, so (by Proposition 6.2 in that paper), and defining $v\left(x\right)\coloneqq\mathbb{E}_{H}u\left(x+W\right)$,
$\begin{split}\bar{\mu}_{u}v\left(\hat{a}_{1}\right)+\left(1-\bar{\mu}_{u}\right)v\left(\hat{a}_{0}\right)&\geq\bar{\mu}_{u}v\left(a_{1}\right)+\left(1-\bar{\mu}_{u}\right)v\left(a_{0}\right)\\ &=\bar{\mu}_{u}v\left(b_{1}\right)+\left(1-\bar{\mu}_{u}\right)v\left(b_{0}\right)\text{,}\end{split}$
where the last line follows from the definition of $\bar{\mu}_{u}$. Consequently, the DM prefers $\hat{a}$ to $b$ for all $\mu\in\left[0,\bar{\mu}_{u}\right]$, strictly so for all $\mu\in\left[0,\bar{\mu}_{u}\right)$.
Case 2. Suppose now that $\bar{\mu}_{u}<\bar{\mu}$, and suppose for the sake of contradiction that
$\bar{\mu}_{u}v\left(\hat{a}_{1}\right)+\left(1-\bar{\mu}_{u}\right)v\left(\hat{a}_{0}\right)<\bar{\mu}_{u}v\left(b_{1}\right)+\left(1-\bar{\mu}_{u}\right)v\left(b_{0}\right)\text{.}$
But, as $a_{0}>\hat{a}_{0}>b_{0}$ and $b_{1}>\hat{a}_{1}>a_{1}$, this implies that $\hat{a}$ is strictly dominated when the DM has utility $u$. We need to show that this cannot happen.
Claim A.2.
$\hat{a}$ is not strictly dominated when the DM has concave utility $u$.
Proof.
Appealing to the Wald-Pearce lemma, suppose there exists a $\rho\in\left(0,1\right)$ such that for all $\theta\in\left\{0,1\right\}$
$\rho\mathbb{E}_{H}u\left(b_{\theta}+W\right)+(1-\rho)\mathbb{E}_{H}u\left(a_{\theta}+W\right)>\mathbb{E}_{H}u\left(\hat{a}_{\theta}+W\right)\text{.}$
Observe that the function $v$ is concave. Accordingly, for all $\theta\in\left\{0,1\right\}$
$\mathbb{E}_{H}u\left(\rho b_{\theta}+(1-\rho)a_{\theta}+W\right)>\rho\mathbb{E}_{H}u\left(b_{\theta}+W\right)+(1-\rho)\mathbb{E}_{H}u\left(a_{\theta}+W\right)\text{.}$
But as $\hat{a}$ is undominated for a risk-neutral DM, for this $\rho$, there exists a $\theta\in\left\{0,1\right\}$ such that
$\begin{split}\mathbb{E}_{H}\left[\hat{a}_{\theta}+W\right]&\geq\rho\mathbb{E}_{H}\left[b_{\theta}+W\right]+(1-\rho)\mathbb{E}_{H}\left[a_{\theta}+W\right]\\ \hat{a}_{\theta}&\geq\rho b_{\theta}+\left(1-\rho\right)a_{\theta}\end{split}\text{.}$
Combining these inequalities, we get that for some $\theta\in\left\{0,1\right\}$,
$\mathbb{E}_{H}u\left(\hat{a}_{\theta}+W\right)>\mathbb{E}_{H}u\left(\hat{a}_{\theta}+W\right)\text{,}$
a contradiction. ∎
Now let $a_{0}<a_{1}$. As before, it suffices to show the result when $\hat{a}$ is a $\bar{\mu}-$MPS of $a$, as subsequent dominance can only improve $\hat{a}$’s attractiveness in relation to $b$. This implies
$b_{0}<\hat{a}_{0}<a_{0}<a_{1}<\hat{a}_{1}<b_{1}\text{.}$
Consequently, the chain of inequalities in Expression A.1 can be replicated, replacing function $u$ with $v$.∎
Lemma A.3.
Suppose the transformation from $a$ to $\hat{a}$ is non-monotone. $\hat{a}$ is $b$- and wealth-superior to $a$ if $a_{0}>a_{1}$, $b_{1}>\hat{a}_{1}$, and Inequality 3.1 holds.
Proof.
This replicates the proof of Lemma 3.5, using Claim A.2’s statement that increased risk-aversion cannot render undominated actions dominated.∎
Finally,
Lemma A.4.
Let $a_{0}=a_{1}$. Then, $\hat{a}$ is $b$- and wealth-superior to $a$ if $b_{1}>\hat{a}_{1}\geq a_{1}$ and Inequality 3.1 holds.
Proof.
Replicate the chain of inequalities in Expression A.1, replacing function $u$ with $v$. ∎
This completes the proof of the theorem. ∎
A.6 Proposition 6.2 Proof
Proof.
Directly,
$\begin{split}\hat{\mu}_{u}=\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)+u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}&=\frac{\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{\hat{a}_{0}-\hat{b}_{0}}}{\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{\hat{a}_{0}-\hat{b}_{0}}+\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &\geq\frac{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}}{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}+\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &=\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(a_{0}-b_{0}\right)\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(b_{1}-a_{1}\right)\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{b}_{1}-\hat{a}_{1}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(a_{1}\right)}=\bar{\mu}_{u}\text{,}\end{split}$
where the first inequality follows from the Three-chord lemma (Theorem 1.16 in Phelps (2009)), the second inequality from Inequality 3.1, and the third inequality from the Three-chord lemma. ∎