### v2

# How to Make an Action Better

Abstract
For two actions in a decision problem, $a$ and $b$, each of which that produces a state-dependent monetary reward, we study how to robustly make action $a$ more attractive. Action $\hat{a}$ improves upon $a$ in this manner if the set of beliefs at which $a$ is preferred to $b$ is a subset of the set of beliefs at which $\hat{a}$ is preferred to $b$, irrespective of the risk-averse agent’s utility function (in money). We provide a full characterization of this relation and discuss applications in politics, bilateral trade and insurance.
1 Introduction
Enter Alice. We know that given a choice between actions $a$ and $b$–the former of which pays out some (monetary) prize amount $\alpha_{S}$ if the weather is sunny and $\alpha_{R}$ if the weather is rainy; and the latter of which, $\beta_{S}$ and $\beta_{R}$ in sun and rain, respectively–she prefers $a$. We do not know, however, Alice’s belief about the probability of sun, nor do we know her utility function (over wealth). Given these data and facts, and under the assumption that her utility function is increasing in wealth, what are the properties of a third action $\hat{a}$, so that she must also prefer it to $b$? What if we further assume that Alice is risk-averse?
The knee-jerk response to these questions is that it is obvious: when Alice’s utility is known only to be within the class of increasing utility functions, it must be that $\hat{a}$ corresponds to a first-order stochastic dominance improvement over $a$. Likewise, when Alice is known to be risk averse, it must be that $\hat{a}$ corresponds to a second-order stochastic dominance improvement over $a$. However, these answers do not stand up to scrutiny–in particular, we do not know the lotteries produced by Alice’s subjective belief, so we cannot speak directly of dominance of lotteries.
In our two main results, Theorems 3.1 and 3.7, we fully characterize these relations in terms of the primitives of the environment–the state-dependent payoffs to actions $a$, $\hat{a}$, and $b$. This characterization is distribution-free and almost ordinal–the lone cardinality specification needed merely concerns the expected prize amounts of $a$ and $\hat{a}$. More specifically, if $\hat{a}$ becomes more attractive relative to $b$ than $a$, it must have a weakly higher expected monetary payoff when evaluated at the beliefs of indifference between $a$ and $b$.
Theorem 3.1 concerns the case in which Alice is assumed to be risk averse. We show that Alice preferring $a$ to $b$ must imply that she also prefers $\hat{a}$ to $b$ if and only if $\hat{a}$ dominates an action whose payoffs are a convex combination of the payoffs to $a$ and $b$. If we do not assume Alice to be risk averse, our other main result (Theorem 3.7) is essentially a negative one: the set of beliefs for which $\hat{a}$ is preferred to $b$ is a superset of those for which $a$ is preferred to $b$, no matter Alice’s increasing utility function, if and only if $\hat{a}$ dominates $a$ or $b$. Concavity of $u$ already imposes significant structure on what transformations are improving, and convexity does too, but in the opposite manner. Combining these requirements leaves only dominance as necessary (and it is obviously sufficient).
There are are a number of natural extensions to our main research question. First, what if we require robust improvements to an action vis-a-vis not just one but multiple alternatives? In §$\S$4.1 we argue that such robust improvements correspond to dominance improvements except in special cases. Second, what if we ask for robustness to aggregate risk? We reveal in §$\S$4.2 that the conditions for our main theorems remain necessary and sufficient–that is, our definition can be extended–to allow for aggregate risk with no alteration to the results. Third, we discuss briefly modifications to both $a$ and $b$ (§$\S$4.3) and provide a sufficient condition for $\hat{a}$ to be preferred to $b$’s replacement, $\hat{b}$. Fourth, §$\S$7 explores information acquisition. Now the question becomes: how must $a$ relate to $\hat{a}$ in a way that guarantees that $\hat{a}$ is chosen more frequently in Alice’s problem of decision making with flexible information acquisition? With two states, we show that $\hat{a}$ dominating $a$ is necessary and sufficient.
We also discuss several examples. Up first is a political example: in a two-party election, suppose a party is contemplating whether to replace its nominee, candidate $a$ [^1] We agree that this scenario may seem implausible, but indulge us. Given that the opponent is candidate $b$, what are the properties of a candidate that is robustly more appealing versus $b$ than $a$? Why, she must dominate a convex combination of $a$ and $b$ in the sense of our first theorem. Our other two examples posit a status quo trading arrangement–a trade of a risky asset, or a sale of an insurance contract–and asks what contracts must be accepted given that the status quo one is. Again, our main result tells us that they must be convex combinations of sorts (now, with respect to the outside options).
1.1 Related Work
The body of work studying decision-making under uncertainty is sizeable. The work closest to this one is Pease and Whitmeyer (2023). There, we formulate a binary relation between actions: action $a$ is safer than $b$ if the the set of beliefs at which $a$ is preferred to $b$ grows larger, in a set inclusion sense, when we make Alice more risk averse.
Rothschild and Stiglitz (1970) is a seminal work that characterizes (mean-preserving) transformations of lotteries that are preferred by all risk-averse agents. Aumann and Serrano (2008) formulate an “measure of riskiness” of gambles, as do Foster and Hart (2009) (who are subsequently followed up upon by Bali, Cakici, and Chabi-Yo (2011) and Riedel and Hellmann (2015)). Crucially, these indices and measures correspond to inherently stochastic objects–the lotteries at hand. Our conception of an improvement to an action concerns comparisons of state-dependent payoffs, which are themselves non-random objects (they are just real numbers).
Naturally this work is also connected to the broader literature studying actions that are comparatively friendly toward risk. In addition to the aforementioned paper of ours, Pease and Whitmeyer (2023), which, like this one, centers around a decision-maker’s state-dependent payoffs to actions, this literature includes Hammond III (1974), Lambert and Hey (1979), Karlin and Novikoff (1963), Jewitt (1987), and Jewitt (1989). Notably, they are statements about lotteries, viz., random objects.
Ours is vaguely a comparative statics work–we’re changing an aspect of a decision problem and seeing how it affects a decision-maker’s choice. Our robustness criterion as well as the simplicity of our setting distinguishes our work from the standard pieces, e.g., Milgrom and Shannon (1994), Edlin and Shannon (1998), and Athey (2002). The works involving aggregation (Quah and Strulovici (2012), Choi and Smith (2017), and Kartik, Lee, and Rappoport (2023)) are closer still–as this inherently corresponds to distributional robustness–but none leave as free parameters both the distribution over states and the DM’s utility function, as we do. Special mention is due to Curello and Sinander (2019), who conduct a robust comparative statics exercise in which an analyst, with limited knowledge of an agent’s preferences, predicts the agent’s choice across menus.
In our final section, we explore properties of the new action, $\hat{a}$, that make it more likely to be selected than $a$ when the alternative is $b$ if information is endogenously acquired. This property is similar in spirit to the observation of Matějka and McKay (2015) that new actions added to a menu may “activate” previously un-chosen actions. Muller-Itten, Armenter, and Stangebye (2021) provide a full characterization of this phenomenon. One crucial distinction between our analysis and theirs is that they explore additions whereas our modification is a replacement.
2 Model
There is a topological space of states, $\Theta$, which is endowed with the Borel $\sigma$-algebra, and which we assume to be compact and metrizable. $\theta$ denotes a generic element of $\Theta$. We denote the set of all Borel probability measures on $\Theta$ by $\Delta\equiv\Delta\left(\Theta\right)$. There is also a decision-maker (DM), who is endowed with two actions. $A$ denotes the set of actions, and each action $a\in A$ is a continuous function from the state space to the set of outcomes, $a\colon\Theta\to\mathbb{R}$. Given a probability distribution over states $\mu\in\Delta$, an action is merely a (simple) lottery.
The DM is an expected-utility maximizer, with a von Neumann-Morgenstern utility function defined on the outcome space $u\colon\mathbb{R}\to\mathbb{R}$. We posit that $u$ is strictly increasing, weakly concave, and continuous. We further specify that no action $a\in A$ is weakly dominated by another. On occasion, we will drop the assumption that $u$ is weakly concave, merely requiring that it be strictly increasing (and continuous). For convenience, for any $a\in A$, we write $a_{\theta}\equiv a\left(\theta\right)$.
Given two actions $\left(a,b\right)$, we are interested in transformations of $a$ that affect the DM’s choice in a particular way. That is, we will modify $a$ to some new $\hat{a}$, in which case the new pair is $\left(\hat{a},b\right)$. To that end, we define the set $P_{b}\left(a\right)$ to be the subset of the probability simplex on which action $a$ is weakly preferred to $b$; formally,
$P_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\right\}\text{.}$
$\bar{P}_{b}\left(a\right)$, in turn, is the subset of the probability simplex on which $a$ is strictly preferred to $b$:
$\bar{P}_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\right\}\text{.}$
When $a$ is transformed to $\hat{a}$, we define the analogous sets $P_{b}\left(\hat{a}\right)$ and $\bar{P}_{b}\left(\hat{a}\right)$.
Definition 2.1.
Action $\hat{a}$ is $b$-Superior to action $a$ if, no matter the risk-averse agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{b}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{b}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $b$.
To put differently, action $\hat{a}$ is $b$-superior to action $a$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, concave, and continuous $u$.
3 Robust Improvements
We partition $\Theta$ into three sets:
$\begin{split}\mathcal{A}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>b_{\theta}\right\}\quad\mathcal{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=b_{\theta}\right\}\text{.}\end{split}$
Given actions $a$ and $b$ and $\lambda\in\left[0,1\right]$, we define the Mixture of $a$ and $b$ to be the action $a^{\lambda}$ that yields payoff
$a^{\lambda}_{\theta}\coloneqq\lambda a_{\theta}+\left(1-\lambda\right)b_{\theta}$
in each state $\theta\in\Theta$. Our main result is, then,
Theorem 3.1.
Fix $a$ and $b$. Action $\hat{a}$ is $b$-superior to action $a$ if and only if $\hat{a}_{\theta}>b_{\theta}$ for all $\theta\in\mathcal{A}$ and $\hat{a}$ dominates a mixture of $a$ and $b$.
Proof.
[…]
Suppose $\mu\in\Delta$ is such that the DM prefers $a$ to $b$, i.e., $L_{a}\succeq L_{b}$. By independence (which is implied by expected utility), for any $\lambda\in\left[0,1\right]$,
$L^{\lambda}\coloneqq\lambda L_{a}+\left(1-\lambda\right)L_{b}\succeq L_{b}\text{.}$
Moreover, by assumption there is some $\lambda^{*}\in\left[0,1\right]$ for which $\lambda^{*}a_{\theta}+\left(1-\lambda^{*}\right)b_{\theta}=\hat{a}_{\theta}$ for all $\theta\in\Theta$. This implies that $L^{\lambda^{*}}$ is a mean preserving spread of $L_{\hat{a}}$. Consequently, as the DM is risk averse, $L_{\hat{a}}\succeq L^{\lambda^{*}}\succeq L_{b}$.∎
This sufficiently result holds for a DM with preferences other than expected utility: we say a preference $\succ$ over lotteries satisfies Betweenness (Hong (1983), Dekel (1986)) if
$P\succ Q\quad\Rightarrow\quad P\succ\alpha P+(1-\alpha)Q\succ Q\text{ for all }\alpha\in\left(0,1\right)\text{.}$
A preference over lotteries is Strongly risk averse (Maccheroni, Marinacci, Wang, and Wu (2023)) if
$Q\text{ is a mean-preserving spread of }P\quad\Rightarrow\quad P\succeq Q\text{.}$
Corollary 3.2.
Suppose the DM’s preferences satisfy betweenness and strong risk aversion. Then, if $\hat{a}_{\theta}>b_{\theta}$ for all $\theta\in\mathcal{A}$ and $\hat{a}$ dominates a mixture of $a$ and $b$, $\hat{a}$ is $b$-superior to action $a$.
Necessity is a little more involved. We prove it in two steps. We say that $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$ if for any pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$, there exists a $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta^{\prime}}\text{.}$
Then, the first of the two necessity steps is
Lemma 3.3.
If $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$ for all $\theta^{\dagger}\in\mathcal{C}$, and $\hat{a}_{\theta}>b_{\theta}$ for all $\theta\in\mathcal{A}$, then $\hat{a}$ dominates a mixture of $a$ and $b$.
Proof.
Suppose $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$ for all $\theta^{\dagger}\in\mathcal{C}$, and $\hat{a}_{\theta}>b_{\theta}$ for all $\theta\in\mathcal{A}$. This implies that for any pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$,
$\min\left\{1,\frac{\hat{a}_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}}\right\}\geq\lambda_{\theta,\theta^{\prime}}\geq\max\left\{0,\frac{b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}{b_{\theta^{\prime}}-a_{\theta^{\prime}}}\right\}\text{.}$
This implies that there exists a $\lambda\in\left[0,1\right]$ for which
$\inf_{\theta\in\mathcal{A}}\left\{1,\frac{\hat{a}_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}}\right\}\geq\lambda\geq\sup_{\theta^{\prime}\in\mathcal{B}}\left\{0,\frac{b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}{b_{\theta^{\prime}}-a_{\theta^{\prime}}}\right\}\text{,}$
and so $\hat{a}$ dominates a mixture of $a$ and $b$.∎
We have shown that pairwise-dominance implies the existence of a common convex weight. That brings us to the second and final necessity step, which closes the equivalence circle stated in the theorem. Namely, we show that pairwise dominance is necessary.
Lemma 3.4.
$\hat{a}$ is $b$-superior to $a$ only if i. for any $\theta^{\dagger}\in\mathcal{C}$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$; ii. for any $\theta\in\mathcal{A}$, $\hat{a}_{\theta}>b_{\theta}$; and iii. $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$.
Proof.
The first two conditions are obviously necessary for $b$-superiority. We assume they hold and suppose for the sake of contraposition that $\hat{a}$ does not pairwise-dominate a collection of mixtures of $a$ and $b$. This means that there exists at least one pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$ for which there exists no $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that the two inequalities in Expression 3 hold.
Fix such a pair of $\theta$ and $\theta^{\prime}$. Given utility $u$, let $\bar{\mu}_{u}$ solve
$\left(1-\bar{\mu}_{u}\right)u\left(b_{\theta}\right)+\bar{\mu}_{u}u\left(b_{\theta^{\prime}}\right)=\left(1-\bar{\mu}_{u}\right)u\left(a_{\theta}\right)+\bar{\mu}_{u}u\left(a_{\theta^{\prime}}\right)\text{,}$
and $\hat{\mu}_{u}$ solve
$\left(1-\hat{\mu}_{u}\right)u\left(b_{\theta}\right)+\hat{\mu}_{u}u\left(b_{\theta^{\prime}}\right)=\left(1-\hat{\mu}_{u}\right)u\left(\hat{a}_{\theta}\right)+\hat{\mu}_{u}u\left(\hat{a}_{\theta^{\prime}}\right)\text{;}$
viz., these are the beliefs on the edge of $\Delta$ between $\delta_{\theta}$ and $\delta_{\theta^{\prime}}$ at which the DM with utility $u$ is indifferent between $a$ and $b$ and $\hat{a}$ and $b$, respectively. If the DM is risk-neutral, we drop the subscript and simply write $\bar{\mu}$ and $\hat{\mu}$.
By assumption $\bar{\mu},\hat{\mu}\in\left(0,1\right)$. Moreover, observe that a necessary condition for $b$-superiority is $\hat{\mu}\geq\bar{\mu}$, which holds if and only if
$\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)\left(\hat{a}_{\theta}-b_{\theta}\right)\geq\left(a_{\theta}-b_{\theta}\right)\left(b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}\right)\text{.}$
If this inequality does not hold, we are done, as we have produced the right negation. Accordingly, suppose it does hold. This means, given our other assumptions,
$b_{\theta}<a_{\theta}<\hat{a}_{\theta},\quad\text{and}\quad\hat{a}_{\theta^{\prime}}<a_{\theta^{\prime}}<b_{\theta^{\prime}}\text{.}$
Our goal is to show that $\bar{\mu}_{u}>\hat{\mu}_{u}$ for some strictly-increasing, continuous, and concave $u$. Let us specify a particular such $u$. Let
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{\theta^{\prime}},\\ \iota x+(1-\iota)a_{\theta^{\prime}},\quad&\text{if}\quad x>a_{\theta^{\prime}}\text{.}\end{cases}$
for some $\iota\in(0,1)$, so $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)}{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)+\iota\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)}-\frac{u\left(\hat{a}_{\theta}\right)-u\left(b_{\theta}\right)}{u\left(\hat{a}_{\theta}\right)-u\left(b_{\theta}\right)+\iota\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)+a_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}\text{.}$
If $b_{\theta}\geq a_{\theta^{\prime}}$, this equals
$\frac{a_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}+b_{\theta^{\prime}}-a_{\theta^{\prime}}}-0>0\text{,}$
when $\iota=0$. If $b_{\theta}<a_{\theta^{\prime}}$, this equals
$1-\frac{u\left(\hat{a}_{\theta}\right)-b_{\theta}}{u\left(\hat{a}_{\theta}\right)-b_{\theta}+a_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}>0\text{,}$
when $\iota=0$.∎
3.1 Intuition
To gain intuition for the theorem, we now discuss the two-state environment. The state is either 00 or $1$, and we specify without loss of generality that $a_{0}>b_{0}$ and $b_{1}>a_{1}$. We let $\mu\in\left[0,1\right]$ denote the probability that the state is $1$.
Here, then, is a purely geometric perspective on Theorem 3.1. Take the line, in belief $\mu$,
$\ell_{a}(\mu)=a_{0}(1-\mu)+a_{1}\mu\text{,}$
which is the expected payoff (in money) of action $a$ as a function of the belief $\mu\in\left[0,1\right]$. In turn, lines
$\ell_{\hat{a}}(\mu)=\hat{a}_{0}(1-\mu)+\hat{a}_{1}\mu\quad\text{and}\quad\ell_{b}(\mu)=b_{0}(1-\mu)+b\mu\text{,}$
are the expected payoffs to actions $\hat{a}$ and $b$, respectively.
Recalling that $\bar{\mu}$ is the value of $\mu$ at which the expected (monetary) payoffs to actions $a$ and $b$ are the same;
$\bar{\mu}=\frac{a_{0}-b_{0}}{a_{0}-b_{0}+b_{1}-a_{1}}\text{,}$
$b$-superiority of $\hat{a}$ is equivalent to $\ell_{\hat{a}}$ lying weakly above a counter-clockwise rotation of $\ell_{a}$ around the point $\left(\bar{\mu},\ell_{a}\left(\bar{\mu}\right)\right)$, provided the rotation isn’t too extreme so that $\hat{a}_{0}\leq b_{0}$. When $\ell_{\hat{a}}$ is a counter-clockwise rotation of $\ell_{a}$, we have made it “more like $\ell_{b}$,” and this cannot hurt $\hat{a}$’s comparative attractiveness versus $b$.
Figure 1 illustrates the benefits of a counter-clockwise rotation. The top three lines are the expected payoffs, in belief $\mu$, to $b$, $\hat{a}$, and $a$; in dashed red, dotted green, and solid blue, respectively. The bottom three lines are the expected utilities (with $u\left(\cdot\right)=\sqrt{1+\cdot}-3$) to the three actions, maintaining the same color and fill correspondences. Line $x=\bar{\mu}_{u}$ is in (thin) solid black and $x=\hat{\mu}_{u}$ is in (thin) dotted orange. Here is an interactive link, where the line can be rotated by moving the $l$ slider. See how $\hat{\mu}_{u}$ moves!
Figure 1: Improving $a$. Try it yourself (rotate by moving the slider)!
Before going on, let us jot down this informal discussion as a formal result.
Corollary 3.5.
$\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and there exists a $\lambda\in\left[0,1\right]$ such that $\ell_{\hat{a}}$ lies weakly above the line $\lambda\ell_{a}+\left(1-\lambda\right)\ell_{b}$.
3.2 A Broader Class of Utilities
So far we have assumed mild but meaningful structure on the DM’s utility function. More is better ($u$ is strictly increasing in the reward), and the DM is risk-averse ($u$ is, at least weakly, concave). Suppose we remove the assumption of risk aversion. We still assume that $u$ is strictly increasing but now allow it to take arbitrary shape.
As we now document, this is, in a sense, a step too far. We will show that without the assumption of risk aversion, the only transformations that must make $b$ less attractive are unambiguous dominance improvements from $a$ to $\hat{a}$. This is intuitive; think back to Corollary 3.5. There, the necessary relationship between $\hat{a}$ and $a$ is $\ell_{\hat{a}}$’s dominance of a counterclockwise rotation of $\ell_{a}$. However, the opposite modification is needed when $u$ is convex–which is now permitted–that is, $\ell_{\hat{a}}$ needs to dominate a clockwise rotation of $\ell_{a}$. This leaves only dominance improvements.
Definition 3.6.
Action $\hat{a}$ is $b$-Better than action $a$ if, no matter the (not necessarily risk-averse) agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{b}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{b}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $b$
To put differently, action $\hat{a}$ is $b$-better than action $a$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, but not necessarily concave, continuous $u$.
Theorem 3.7.
Action $\hat{a}$ is $b$-better than action $a$ if and only if for all $\theta\in\mathcal{A}$ $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ dominates $a$ or $b$.
Proof.
[…]
$b_{\theta}<\hat{a}_{\theta}<a_{\theta},\quad\text{and}\quad a_{\theta^{\prime}}<\hat{a}_{\theta^{\prime}}<b_{\theta^{\prime}}\text{.}$
Our goal is to show that $\bar{\mu}_{u}>\hat{\mu}_{u}$ for some strictly-increasing, continuous $u$. Take
$u(x)=\begin{cases}\iota x+(1-\iota)\hat{a}_{0},\quad&\text{if}\quad x\leq\hat{a}_{\theta},\\ x,\quad&\text{if}\quad x>\hat{a}_{\theta}\text{,}\end{cases}$
for some $\iota\in(0,1)$. Thus, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{\theta}-\hat{a}_{\theta}-\iota\left(b_{\theta}-\hat{a}_{\theta}\right)}{a_{\theta}-\hat{a}_{\theta}-\iota\left(b_{\theta}-\hat{a}_{\theta}\right)+u(b_{\theta^{\prime}})-u(a_{\theta^{\prime}})}-\frac{\iota\left(\hat{a}_{\theta}-b_{\theta}\right)}{\iota\left(\hat{a}_{\theta}-b_{\theta}\right)+u(b_{\theta^{\prime}})-u(\hat{a}_{\theta^{\prime}})}\text{.}$
If $b_{\theta^{\prime}}\leq\hat{a}_{\theta}$, this equals
$1-\frac{\hat{a}_{\theta}-b_{\theta}}{\hat{a}_{\theta}-b_{\theta}+b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}>0\text{,}$
when $\iota=0$. If $b_{\theta^{\prime}}>\hat{a}_{\theta}$, it equals
$\frac{a_{\theta}-\hat{a}_{\theta}}{a_{\theta}-\hat{a}_{\theta}+b_{\theta^{\prime}}-u(a_{\theta^{\prime}})}-0>0\text{,}$
when $\iota=0$. ∎
4 Extensions
4.1 One Versus Many
What if we look for improvements to $a$ that make it more attractive in comparison to multiple alternatives, not just a single one? For simplicity, we restrict our attention to the two-state environment, $\Theta=\left\{0,1\right\}$. Let $A$ be the DM’s finite set of actions, where we assume that none are (strictly) dominated. Indexing the actions by $I=\left\{1,\dots,m\right\}$ (with $m\geq 3$), we label the actions so that for all $i\in I$, $a_{0}^{i}>a_{0}^{j}$ and $a_{1}^{i}<a_{1}^{j}$ for all $j>i$ ($j\in I$).
There are two special actions, $a^{1}$ and $a^{m}$. The first is uniquely optimal in state 00 and the second is uniquely optimal in state $1$. We term these the two extreme actions. For a specified $a\in A$, we define $B\coloneqq A\setminus\left\{a\right\}$ and
$P_{B}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\max_{a^{j}\in B}\mathbb{E}_{\mu}u\left(a^{j}_{\theta}\right)\right\}\text{.}$
We define $\bar{P}_{B}\left(a\right)$ to be the analogous set with a strict inequality, and sets $P_{B}\left(\hat{a}\right)$ and $\bar{P}_{B}\left(\hat{a}\right)$ their compatriots once we’ve replaced $a$ with $\hat{a}$.
Definition 4.1.
Action $\hat{a}$ is $B$-Superior to action $a$ if the set of beliefs at which the DM prefers $a$ to the actions other than $a$, $P_{B}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to those other actions, $P_{B}(\hat{a})$; and $\bar{P}_{B}(a)\subseteq\bar{P}_{B}(\hat{a})$; no matter the risk-averse agent’s utility function.
We have the following stark result.
Proposition 4.2.
If $a=a^{1}$ ($a=a^{m}$), $\hat{a}$ is $B$-superior to $a$ if and only if $a$ is $b$-superior to $a^{2}$ ($a^{m-1}$). If $a$ is not one of the extreme actions, $\hat{a}$ is $B$-superior to $a$ if and only if $\hat{a}$ dominates $a$.
Proof.
[…]
If $a\neq a^{1},a^{m}$ it is immediate that $\hat{a}$ dominating $a$ implies $P_{B}(a)\subseteq P_{B}(\hat{a})$.
[…]
The logic behind this proposition is straightforward. For an action to improve versus the action to its right, we need the induced (risk-neutral) line in belief space to lie above some counterclockwise rotation around the indifference point between the two actions. For an action to improve versus the action to its left, we need the line in belief space to lie above some clockwise rotation about the indifference point between the two actions (which is strictly to the left of the other indifference point). A dominance shift is the only way to reconcile these requirements.
4.2 Robustness to Aggregate Risk
What if the DM also has unknown (to us) random wealth? Formally, in any state $\theta$, the DM’s utility from action $a$ is $\mathbb{E}_{H}u\left(a_{\theta}+W\right)$, where $W$ is a state-independent (finite-mean) random variable .
Abusing notation, we define $P_{a,b}(a)$ to be the subset of $\Delta$ for which $a$ is preferred to $b$:
$P_{b}\left(a\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu,H}u\left(a_{\theta}+W\right)\geq\mathbb{E}_{\mu,H}u\left(b_{\theta}+W\right)\right\}\text{.}$
$\bar{P}_{b}\left(a\right)$ is this set’s strict-preference counterpart.
We extend our definitions in the natural way:
Definition 4.3.
Action $\hat{a}$ is $b$- and Wealth-Superior to action $a$ if the set of beliefs at which the DM prefers $a$ to $b$, $P_{b}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to $b$, $P_{b}(\hat{a})$; and the set of beliefs at which the DM strictly prefers $a$ to $b$, $\bar{P}_{b}(a)$, is a subset of the set of beliefs at which the DM strictly prefers $\hat{a}$ to $b$, $\bar{P}_{b}(\hat{a})$; no matter the risk-averse agent’s utility function (for any strictly increasing, concave $u$) and no matter the agent’s state-independent random wealth $W$ (for any finite-mean, real-valued random variable $W$).
Definition 4.4.
Action $\hat{a}$ is $b$- and Wealth-Better than action $a$ if the set of beliefs at which the DM prefers $a$ to $b$, $P_{b}(a)$, is a subset of the set of beliefs at which the DM prefers $\hat{a}$ to $b$, $P_{b}(\hat{a})$; and the set of beliefs at which the DM strictly prefers $a$ to $b$, $\bar{P}_{b}(a)$, is a subset of the set of beliefs at which the DM strictly prefers $\hat{a}$ to $b$, $\bar{P}_{b}(\hat{a})$; no matter the (not necessarily risk-averse) agent’s utility function (for any strictly increasing, concave $u$) and no matter the agent’s state-independent random wealth $W$ (for any finite-mean, real-valued random variable $W$).
Then,
Theorem 4.5.
$\hat{a}$ is $b$- and wealth-superior to $a$ if and only if $\hat{a}$ is an improvement to $b$. $\hat{a}$ is $b$- and wealth-better than $a$ if and only if for all $\theta\in\mathcal{A}$ $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ dominates $a$ or $b$.
Proof.
[…]
Let $Y_{a}$ be the random variable yielding $a_{\theta}$ with probability $\mu\left(\theta\right)$, and $L_{a}$ the sum of $Y_{a}$ and $W$; $Y_{b}$, the random variable yielding $b_{\theta}$ with probability $\mu\left(\theta\right)$, and $L_{b}$ the sum of $Y_{b}$ and $W$; and $Y_{\hat{a}}$, the random variable yielding $\hat{a}_{\theta}$ with probability $\mu\left(\theta\right)$, and $L_{\hat{a}}$ the sum of $Y_{\hat{a}}$ and $W$.
Suppose $\mu\in\Delta$ and $W$ are such that the DM prefers $a$ to $b$, i.e., $L_{a}\succeq L_{b}$. By independence, for any $\lambda\in\left[0,1\right]$,
$L^{\lambda}\coloneqq\lambda L_{a}+\left(1-\lambda\right)L_{b}\succeq L_{b}\text{.}$
Moreover, by assumption there is some $\lambda^{*}\in\left[0,1\right]$ for which $\lambda^{*}a_{\theta}+\left(1-\lambda^{*}\right)b_{\theta}=\hat{a}_{\theta}$ for all $\theta\in\Theta$. This implies that $\lambda^{*}Y_{a}+\left(1-\lambda^{*}\right)Y_{b}$ is a mean preserving spread of $Y_{\hat{a}}$. Moreover, as second-order stochastic dominance is preserved under independent noise, [^3] See, for instance, the discussion on page 3 of Pomatto, Strack, and Tamuz (2020). $L^{\lambda^{*}}=\lambda^{*}Y_{a}+\left(1-\lambda^{*}\right)Y_{b}+W$ is a mean preserving spread of $L_{\hat{a}}=Y_{\hat{a}}+W$. Consequently, as the DM is risk averse, $L_{\hat{a}}\succeq L^{\lambda^{*}}\succeq L_{b}$. ∎
4.3 Everything Everywhere All at Once
In this paper, we focus on altering a single action, leaving the other action (or actions) unaltered. This is reasonable in many situations–for example, in the two applications in the next section, the other action is the DM’s outside option, which is unalterable. But what if we change both actions $a$ and $b$? As we will see, the basic intuition from the earlier results remains true, though we only provide a partial analysis, primarily to avoid getting bogged down in a tedium of cases. We also restrict attention to the two-state environment ($\Theta=\left\{0,1\right\}$), as it is straightforward to aggregate the conditions (as in Pease and Whitmeyer (2023)) to extend the result to a general state space.
Now, we are going from $a$ to $\hat{a}$ and $b$ to some $\hat{b}$. To that end, we define
$P_{\hat{b}}\left(\hat{a}\right)\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(\hat{b}_{\theta}\right)\right\}\text{,}$
understanding $\bar{P}_{\hat{b}}\left(\hat{a}\right)$ to be this set’s strict inequality counterpart.
Definition 4.6.
Given $a$ and $b$, action $\hat{a}$ is $\hat{b}$-Superior to action $a$ if, no matter the risk-averse agent’s utility function,
  1. 1.
    $P_{b}(a)\subseteq P_{\hat{b}}(\hat{a})$ and
  2. 2.
    $\bar{P}_{b}(a)\subseteq\bar{P}_{\hat{b}}(\hat{a})$
i.e., the set of beliefs at which the DM prefers $a$ to $b$ is a subset of the beliefs at which the DM prefers $\hat{a}$ to $\hat{b}$.
The following inequality–which is the analog of Inequality 3 when we make a double-modification–is necessary for $\hat{b}$-superiority when $\hat{a}_{1}<\hat{b}_{1}$:
$\left(b_{1}-a_{1}\right)\left(\hat{a}_{0}-\hat{b}_{0}\right)\geq\left(a_{0}-b_{0}\right)\left(\hat{b}_{1}-\hat{a}_{1}\right)\text{.}$
Obviously, $\hat{a}$ dominating $a$ or $\hat{b}$ and $\hat{b}$ being dominated by $b$ or $\hat{a}$ imply $\hat{b}$-superiority. Here is a less obvious sufficient condition.
Proposition 4.7.
If Inequality 4.3 holds, $a_{1}\leq\hat{a}_{1}<b_{1}\leq\hat{b}_{1}$, and $\hat{b}_{0}\leq b_{0}<\hat{a}_{0}\leq a_{0}$, $\hat{a}$ is $\hat{b}$-superior to action $a$.
Proof.
The proof is purely mechanical, so we relegate it to Appendix A.1. ∎
As we discussed above, $\hat{a}$ improving upon $a$ versus $b$ (when $b$ is fixed) is equivalent to $\hat{a}$’s expected payoff in belief space lying above a counter-clockwise rotation of $a$’s payoff. When we alter both actions, the rotation point may now shift–think vaguely of a co-movement of payoffs to both $a$ and $b$–and $\hat{a}$’s payoff and $\hat{b}$’s payoffs are both counter-clockwise rotations about this point.
4.4 A Menu of Three
Our conception of superiority does not mean that $\hat{a}$ must be preferred to $a$ given the DM’s utility function and belief, only to $b$. This begs the following question. What conditions are equivalent to the DM choosing $\hat{a}$ over both $a$ and $b$?
We say that action $\hat{a}$ is $\left(a,b\right)$-superior if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\max_{c\in\left\{a,b\right\}}\mathbb{E}_{\mu}u\left(c_{\theta}\right)\text{,}$
for any strictly increasing, concave, and continuous $u$. Then,
Proposition 4.8.
$\hat{a}$ is $\left(a,b\right)$-superior if and only if $\hat{a}$ dominates a mixture of $a$ and $b$ and $\hat{a}_{\theta}\geq a_{\theta}$ for all $\theta\in\mathcal{A}$.
Proof.
[…]
$\mathbb{E}_{\mu^{\prime}}u\left(\hat{a}_{\theta}\right)\geq\max_{c\in\left\{a,b\right\}}\mathbb{E}_{\mu^{\prime}}u\left(c_{\theta}\right)\text{.}$
By the linearity of the expectation, for any $\mu$ in the closed convex hull of $\operatorname*{extreme}P_{b}(a)$, the DM prefers $\hat{a}$ to both $a$ and $b$. By the Krein-Milman theorem, this set is precisely $P_{b}(a)$, as desired.
[…]
4.5 A Hidden Alternative
Suppose $b$ is not known; or rather, the DM’s state-dependent payoffs to $b$ are not known. It is possible, nevertheless, to say something as follows.
Corollary 4.9.
Suppose action $b$ is dominated by some $\tilde{b}$. Then, $\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}$ dominates a mixture of $a$ and $\tilde{b}$.
We omit the proof as it is an easy consequence of Theorem 3.1. If we think of $b$ as the DM’s “outside option,” this corollary tells us that we can formulate a meaningful robust alteration of action $a$ even when the outside option is not fully specified. All we need is an “upper bound” of sorts on the outside option.
5 Applications
5.1 Choosing Which Candidate to Run
There is a unit mass of voters and two candidates $a$ and $b$ running for office. There is an unknown state of the world $\theta\in\Theta$, and candidate $a$ ($b$) produces monetary benefit $a_{\theta}$ ($b_{\theta}$) to each voter in every state $\theta\in\Theta$. Neither candidate dominates the other in the sense that there is at least one state of the world in which each provides a strictly higher monetary benefit to the populace than the other.
The party backing candidate $a$ is contemplating whether to replace him with a different candidate $\hat{a}$. Crucially, it knows neither the risk preferences of the populace (which may be heterogeneous) nor the beliefs of the populace (also possibly heterogeneous). Accordingly, it makes its choice based on a robust criterion, asking what are the candidates who have a better chance of winning versus $b$, no matter the population’s (concave) $u$s or beliefs. Who cannot be worse than $a$ versus $b$? Theorem 3.1 tells us the answer:
Remark 5.1.
Candidate $\hat{a}$ is no worse than $a$ versus $b$ if and only if she dominates a mixture of $a$ and $b$.
5.2 Robustly Optimal Bilateral Trade Modifications
There is a buyer ($B$) and a seller ($S$). $S$ possesses an asset that pays out $v_{\theta}\in\mathbb{R}$ in state $\theta$. The state is contractible and the status quo trade agreement sees transfers of size $\gamma_{\theta}\in\mathbb{R}$ from $B$ to $S$ in each state $\theta\in\Theta$. We assume that there exist states $\theta^{\prime},\theta^{\prime\prime}\in\Theta$ for which $v_{\theta^{\prime}}>\gamma_{\theta^{\prime}}$ and $\gamma_{\theta^{\prime\prime}}>v_{\theta^{\prime\prime}}$.
$B$’s state-dependent payoff from transacting is $v_{\theta}-\gamma_{\theta}$, and $S$’s is $\gamma_{\theta}$ (budget balance). $B$’s outside option is the sure-thing 00 and $S$’s outside option is the asset, with random payoff $v_{\theta}$ in each state $\theta\in\Theta$.
We suppose that the status-quo is acceptable–given the agents’ subjective beliefs about uncertainty and their (concave) utility functions, each is willing to participate in the arrangement. What modifications to the arrangement are robustly optimal in that $B$ and $S$ will still remain willing to participate? Theorem 3.1 reveals the answer.
We observe that for all $\theta\in\Theta$ and for any $\lambda\in\left[0,1\right]$,
$\lambda\left(v_{\theta}-\gamma_{\theta}\right)=v_{\theta}-\lambda\gamma_{\theta}-\left(1-\lambda\right)v_{\theta}\text{,}$
so that any new transfer that yields the requisite convex combination of $B$’s status quo payoff from accepting the terms and $B$’s outside option of 00 equals
$\lambda\gamma_{\theta}+\left(1-\lambda\right)v_{\theta}\text{,}$
which is precisely the convex combination (with the same weight) of $S$’s status quo payoff and her outside option. This, plus our assumption of budget balance, imply
Remark 5.2.
A new trade agreement, $\left(\hat{\gamma}_{\theta}\right)_{\theta\in\Theta}$ must be acceptable if and only if there exists some $\lambda\in\left(0,1\right]$ such that $\hat{\gamma}_{\theta}=\lambda\gamma_{\theta}$ for all $\theta\in\Theta$.
5.3 Robustly Optimal Insurance Modifications
Now consider the scenario of a risk-neutral insurer and a consumer. The consumer’s payoff without insurance is $v_{\theta}$ in each state $\theta\in\Theta$. The status-quo policy pays out (net of the loss) $\alpha_{\theta}$ to the consumer in each state $\theta\in\Theta$. We assume that there exist some $\theta^{\prime},\theta^{\prime\prime}\in\Theta$ for which $\alpha_{\theta^{\prime}}>v_{\theta^{\prime}}$ and $\alpha_{\theta^{\prime\prime}}<v_{\theta^{\prime\prime}}$.
Remark 5.3.
A new contract, $\left(\hat{\alpha}_{\theta}\right)_{\theta\in\Theta}$ must be acceptable if and only if there exists some $\lambda\in\left(0,1\right]$ such that $\hat{\alpha}_{\theta}\geq\lambda\alpha_{\theta}+\left(1-\lambda\right)v_{\theta}$ for all $\theta\in\Theta$.
One interesting implication of this remark is that the set of acceptable contracts could very well expose the consumer to more risk. Indeed, suppose that the status-quo policy is risk free, i.e., yields the consumer the same wealth in every state. The consumer’s outside option (not buying insurance) is, of course, riskier. In this case, any actuarily-fair contract that the consumer must accept also exposes her to risk as her wealth is (non-trivially) state-dependent.
6 A New (Partial) Order
We can also construct a novel partial order over actions in a decision problem, the “$b$-superior order.” Fix a state-space $\Theta$, an action $b$, and a set of actions $A$, where $b\in A$ and no action $a\in A\setminus\left\{b\right\}$ dominates $b$. Let $\hat{a}\trianglerighteq a$ denote that action $\hat{a}$ is $b$-superior to action $a$. Then,
Proposition 6.1.
[…]
Proof.
[…]
Antisymmetry. Suppose $a^{1}\trianglerighteq a^{2}$ and $a^{2}\trianglerighteq a^{1}$. Then there exist $\lambda_{1},\lambda_{2}\in\left[0,1\right]$ such that
$a^{1}_{\theta}\geq\lambda_{1}a^{2}_{\theta}+\left(1-\lambda_{1}\right)b_{\theta}\quad\forall\ \theta\in\Theta\text{,}$
and
$a^{2}_{\theta}\geq\lambda_{2}a^{1}_{\theta}+\left(1-\lambda_{2}\right)b_{\theta}\quad\forall\ \theta\in\Theta\text{.}$
Suppose for the sake of contradiction that there exists a $\theta^{\prime}\in\Theta$ for which $a^{1}_{\theta^{\prime}}>a^{2}_{\theta^{\prime}}$. This implies that $\lambda_{2}<1$.
Combining the two inequalities, and doing a little algebra, we get that
$a^{1}_{\theta}\left(1-\lambda_{1}\lambda_{2}\right)\geq\left(1-\lambda_{1}\lambda_{2}\right)b_{\theta}\quad\forall\ \theta\in\Theta\text{,}$
a contradiction as $01-\lambda_{1}\lambda_{2}>$ 0 and $a^{1}$ doesn’t dominate $b$.
Transitivity. Suppose $a^{1}\trianglerighteq a^{2}$ and $a^{2}\trianglerighteq a^{3}$. Then there exist $\lambda_{1},\lambda_{2}\in\left[0,1\right]$ such that Inequality 6 and
$a^{2}_{\theta}\geq\lambda_{2}a^{3}_{\theta}+\left(1-\lambda_{2}\right)b_{\theta}\quad\forall\ \theta\in\Theta$
hold. Combining these, we get
$a^{1}_{\theta}\geq\lambda_{1}\lambda_{2}a^{3}_{\theta}+\left(1-\lambda_{1}\lambda_{2}\right)b_{\theta}\quad\forall\ \theta\in\Theta\text{,}$
so $a^{1}\trianglerighteq a^{3}$, as desired.∎
We have come up with a partial order that is similar in spirit to that formulated by Hart (2011). Crucially, his concerns lotteries; ours, families of lotteries.
7 Information Acquisition
Up until this point we have focused on a simple decision problem in which a DM has a belief over the states of the world and chooses an optimal action based on this belief. It is easy to see that $b$-superiority extends to the scenario in which a DM obtains exogenous information before taking a decision. That is, $\hat{a}$ is $b$-superior to $a$ if and only if for any $\mu\in\operatorname{int}\Delta$ and any Bayes-plausible (martingale) distribution over posteriors $F$, the DM chooses $\hat{a}$ with a higher probability than she chooses $a$ (where, in both cases, $b$ is chosen with the complimentary probability).
There has been a recent explosion of interest in problems with endogenous information acquisition. Suppose we want to know what the properties of $\hat{a}$ are that make it chosen more versus $b$ than $a$ is when the DM flexibly acquires information before taking a choice.
We begin with the binary-state environment ($\Theta=\left\{0,1\right\}$) for tractability. We maintain our assumptions that neither $a$ nor $b$ dominates the other and that $a$ is uniquely optimal in state 00, $b$ in state $1$. Given a prior $\mu_{0}\in\left(0,1\right)$, and defining an agent’s value function, in belief $\mu$, as
$V(\mu)\coloneqq\max_{\tilde{a}\in\left\{a,b\right\}}\mathbb{E}_{\mu}u(\tilde{a}_{\theta})\text{,}$
the DM’s flexible information acquisition problem is
[…]
Any solution $F^{*}$ ($\hat{F}^{*}$) to Program 7 (7) produces an optimal choice probability of action $a$ ($\hat{a}$) by the DM, which we define to be $p_{c,\mu_{0},u}\equiv p$ ($\hat{p}_{c,\mu_{0},u}\equiv\hat{p}$).
Definition 7.1.
$\hat{a}$ is selected more than $a$ if for any Bayes-plausible $D$, prior $\mu_{0}\in\operatorname{int}\Delta$, and strictly increasing, concave $u$; for any optimal $p$, there exists an optimal choice probability $\hat{p}\geq p$.
Proposition 7.2.
$\hat{a}$ is selected more than $a$ if and only if $\hat{a}$ dominates $a$ or $b$.
Proof.
A full proof may be found in Appendix A.2. ∎
Most of the work in proving the proposition is in the necessity portion. On the one hand, if $\hat{a}$ is not $b$-superior to $a$, we can find a $u$ such that $\hat{\mu}_{u}<\bar{\mu}_{u}$. Take such a $u$, in which case we can then always find a cost function such that $a$ is selected with probability $1$ when the menu is $\left\{a,b\right\}$ but when the menu is $\left\{\hat{a},b\right\}$, either the DM learns in a nontrivial manner, and hence selects $\hat{a}$ with probability strictly less than $1$, or does not learn but selects $b$ with probability $1$. On the other hand, even if $\hat{a}$ is $b$-superior to $a$ (but dominates neither), it may be that $\hat{p}<p$, as taking a convex combination of actions operates like scaling up the marginal cost of learning. Thus, one can always find a cost function and a prior that roughly“favors $b$–” i.e., is moderately, but not too close to $1$–so that the DM selects $b$ with probability $1$ when the menu is $\left\{\hat{a},b\right\}$ but learns, and so selects $a$ with strictly positive probability, when the menu is $\left\{a,b\right\}$.
7.1 Three or More States
Figure 2: The insufficiency of dominance for three or more states
Alas, when there are three states, $\hat{a}$ dominating $a$ no longer implies that $\hat{a}$ is selected more than $a$, as we now illustrate. Let $\Theta=\left\{0,1,2\right\}$ with $\mu^{2}\coloneqq\mathbb{P}(2)$ and $\mu^{1}\coloneqq\mathbb{P}(1)$. Let $\hat{\alpha}_{2}=3/2=\alpha_{2}+1/2$, $\alpha_{1}=-1=\hat{\alpha}_{1}$ and $\hat{\alpha}_{0}=\alpha_{0}=0$; and $\beta_{\theta}=0$ for all $\theta\in\Theta$. Evidently, $\hat{a}$ weakly dominates $a$ (and $\hat{a}\neq a$).
We have
$V(\mu^{1},\mu^{2})=\max\left\{\mu^{2}-\mu^{1},0\right\},\quad\text{and}\quad\hat{V}(\mu^{1},\mu^{2})=\max\left\{\frac{3}{2}\mu^{2}-\mu^{1},0\right\}\text{.}$
We then define
$W(\mu^{1},\mu^{2})\coloneqq\max\left\{V,\hat{V}+\frac{1}{4}\mu^{1}-\frac{1}{8}\right\}\text{,}$
then, appealing to Lemma A.1 in Whitmeyer (2023), conclude that there exists a UPS cost such that the four points
$\left\{\left(\frac{2}{25},\frac{1}{5}\right),\left(\frac{12}{25},\frac{1}{5}\right),\left(\frac{1}{50},\frac{53}{200}\right),\left(\frac{27}{50},\frac{1}{200}\right)\right\}$
are the support of optimal learning given value function $W$. Accordingly, we take such a cost function and fix prior $\left(3/20,1/5\right)$ then compute $p=33/40>3/4=\hat{p}$.
Figure 2 illustrates this example on the $2$-simplex. The four colored polygons are the regions in which $V$ lies above the rotated value function $\hat{V}$ and either $a$ is optimal (red) or $b$ is optimal (orange); and the rotated value function $\hat{V}$ lies above $V$ and either $\hat{a}$ is optimal (purple) or $b$ is optimal (blue). The prior is the hollow red dot. The black dots are the support of $F^{*}$, and the orange $x$s are the support of $\hat{F}^{*}$.
Leaving the single-dimensional environment corresponding to the two-state simplex engenders this result. What we are doing is taking a value function that in a sense “twist things,” making it so that the learning in the two problems is not along the same line segment. There is enough freedom then to pick an appropriate prior so that $p>\hat{p}$.
Appendix A Omitted Proofs
A.1 Proposition 4.7 Proof
Proof.
Directly,
$\begin{split}\hat{\mu}_{u}=\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)+u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}&=\frac{\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{\hat{a}_{0}-\hat{b}_{0}}}{\frac{u\left(\hat{a}_{0}\right)-u\left(\hat{b}_{0}\right)}{\hat{a}_{0}-\hat{b}_{0}}+\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &\geq\frac{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}}{\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{a_{0}-b_{0}}+\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &=\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(a_{0}-b_{0}\right)\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{a}_{0}-\hat{b}_{0}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+\left(b_{1}-a_{1}\right)\frac{u\left(\hat{b}_{1}\right)-u\left(\hat{a}_{1}\right)}{\hat{b}_{1}-\hat{a}_{1}}}\\ &\geq\frac{u\left(a_{0}\right)-u\left(b_{0}\right)}{u\left(a_{0}\right)-u\left(b_{0}\right)+u\left(b_{1}\right)-u\left(a_{1}\right)}=\bar{\mu}_{u}\text{,}\end{split}$
where the first and third inequalities follow from the Three-chord lemma (Theorem 1.16 in Phelps (2009)), and the second inequality from Inequality 4.3. ∎
A.2 Proposition 7.2 Proof
Proof.
[…]
Suppose first that $\mu_{0}$ and $c$ are such that $p=1$. This is equivalent to $\mu_{0}\leq\bar{\mu}_{u}$ and
$\left(\alpha_{1}-\alpha_{0}-c^{\prime}(\mu_{0})\right)\mu+c^{\prime}(\mu_{0})+\alpha_{0}\geq 0\text{,}$
for all $\mu\in\left[0,1\right]$. For any $\mu\in\left[0,1\right]$ this expression is strictly increasing in both $\alpha_{0}$ and $\alpha_{1}$, so we have
$\left(\hat{\alpha}_{1}-\hat{\alpha}_{0}-c^{\prime}(\mu_{0})\right)\mu+c^{\prime}(\mu_{0})+\hat{\alpha}_{0}\geq 0\text{,}$
for all $\mu\in\left[0,1\right]$, i.e., $\hat{p}=1$.
Now suppose that $\mu_{0}$ is such that $p\in\left(0,1\right)$. In this case, the support of $F^{*}$, $x_{L}$ and $x_{H}$, solves
$\alpha_{1}-\alpha_{0}-c^{\prime}(x_{L})+c^{\prime}(x_{H})=0\text{,}$
and
$c^{\prime}(x_{L})x_{L}-c^{\prime}(x_{H})x_{H}+c(x_{H})+\alpha_{0}-c(x_{L})=0\text{.}$
Appealing to the implicit function theorem, we obtain
$x_{L}^{\prime}(\alpha_{1})=\frac{x_{H}}{c^{\prime\prime}(x_{L})\left(x_{H}-x_{L}\right)}>0,\quad\text{and}\quad x_{H}^{\prime}(\alpha_{1})=\frac{x_{L}}{c^{\prime\prime}(x_{H})\left(x_{H}-x_{L}\right)}>0\text{.}$
Likewise,
$x_{L}^{\prime}(\alpha_{1})=\frac{1-x_{H}}{c^{\prime\prime}(x_{L})\left(x_{H}-x_{L}\right)}>0,\quad\text{and}\quad x_{H}^{\prime}(\alpha_{1})=\frac{1-x_{L}}{c^{\prime\prime}(x_{H})\left(x_{H}-x_{L}\right)}>0\text{.}$
Evidently, $p$ is strictly increasing in both $x_{H}$ and $x_{L}$, so $\hat{p}\geq p$.
[…]
Case iii is immediate. Let the DM be risk neutral: then there exists a $\mu^{\prime}\in\left(0,1\right)$, such that for any belief $\mu<\mu^{\prime}$, it is uniquely optimal for the DM to take action $a$ when her menu is $\left\{a,b\right\}$ and $b$ when her menu is $\left\{\hat{a},b\right\}$. Then, one need only pick a sufficiently convex $c$–that such a convex $c$ can always be found is an implication of Lemma A.1 in Whitmeyer (2023) [^5] This is not strictly true as Whitmeyer (2023) does not impose that the cost function is twice continuously differentiable, merely strictly convex. However, it is easy to extend that result to smooth functions: see, e.g. (2019) (https://mathoverflow.net/users/943/dmitri panov). Alternatively, one could remove the twice-continuously differentiable specification at the expense of not being able to appeal to the implicit function theorem in the sufficiency portion of the proof.–such that for prior $\mu_{0}$, no learning is uniquely optimal in both problems, in which case $p=1>0=\hat{p}$.
Case i is also easy. Observe that $\hat{a}$ is not $b$-superior to $a$, so there exists a strictly increasing concave $u$ for which $10<\hat{\mu}_{u}<\bar{\mu}_{u}<$ 1. We maintain the convention $\alpha_{\theta}\equiv u(a_{\theta})$ and $\hat{\alpha}_{\theta}\equiv u(\hat{a}_{\theta})$ for all $\theta\in\left\{0,1\right\}$ and also introduce the notation $\beta_{\theta}\equiv u(b_{\theta})$ for all $\theta\in\left\{0,1\right\}$.
We tweak the notation
$\ell_{a}=\mu\alpha_{1}+\left(1-\mu\right)\alpha_{0},\ \ell_{\hat{a}}=\mu\hat{\alpha}_{1}+\left(1-\mu\right)\hat{\alpha}_{0},\text{ and }\ell_{b}=\mu\beta_{1}+\left(1-\mu\right)\beta_{0}\text{,}$
and define $W(\mu)\coloneqq\max\left\{\ell_{a},\ell_{\hat{a}},\ell_{b}\right\}$. We let $\tilde{\mu}$ denote the intersection of $\ell_{a}$ and $\ell_{\hat{a}}$, and observe that $0<\tilde{\mu}<\bar{\mu}$; this holds because $\ell_{\hat{a}}$ has a steeper slope and a strictly larger $y$-intercept than $\ell_{a}$. Then, Lemma A.1 in Whitmeyer (2023) implies that for any triple $\mu_{1}\in\left(0,\tilde{\mu}\right)$, $\mu_{2}\in\left(\tilde{\mu},\bar{\mu}\right)$, and $\mu_{3}\in\left(\bar{\mu},1\right)$, there exists a UPS cost such that when the DM’s value function is $W$, any optimal learning has support on the three specified points. Accordingly, for such a cost function, when $\mu_{0}=\mu_{2}$, $p=1$ and $\hat{p}<1$.
Finally, case ii: the argument from the previous paragraph allows us to assume that $\hat{a}$ is $b$-superior to $a$, or else we are done. Consequently, $\bar{\mu}_{u}\leq\hat{\mu}_{u}$ for all permissible $u$. Fix such a $u$ and normalize payoffs so that $\beta_{\theta}=0$ for all $\theta\in\left\{0,1\right\}$ (this is without loss of generality, as $u$ has been fixed). Now take a line $f(\mu)\coloneqq-\gamma\mu+\delta$, where $0<\delta<\alpha_{0}-\hat{\alpha}_{0}$, $\gamma>\delta$, and
$\frac{\delta}{\gamma}>\hat{\mu}_{u}=\frac{\hat{\alpha}_{0}}{\hat{\alpha}_{0}-\hat{\alpha}_{1}}\text{.}$
Defining $T(\mu)\coloneqq\max\left\{\ell_{a},\max\left\{\ell_{\hat{a}},0\right\}+f,0\right\}$, we note that this piecewise-affine curve has three kink points. First, at some $\mu_{1}\in\left(0,1\right)$, where $\ell_{a}$ and $\ell_{\hat{a}}+f$ intersect. Second, at some $\mu_{2}\in\left(\mu_{1},1\right)$, where 00 and $\ell_{\hat{a}}$ intersect. Third, at some $\mu_{3}\in\left(\mu_{2},1\right)$ where $f$ and 00 intersect. Again appealing to Lemma A.1 in Whitmeyer (2023), taking a prior $\mu_{0}\in\left(\mu_{2},\mu_{3}\right)$ we note the existence of a UPS cost producing $p>0$ and $\hat{p}=0$.∎