### v4

# How to Make an Action Better

Abstract
For two actions in a decision problem, $a$ and $b$, each of which produces a state-dependent monetary reward, we study how to robustly make action $a$ more attractive. Action $\hat{a}$ improves upon $a$ in this manner if the set of beliefs at which $a$ is preferred to $b$ is a subset of the set of beliefs at which $\hat{a}$ is preferred to $b$, irrespective of the risk-averse agent’s utility function (in money). We provide a full characterization of this relation and discuss applications in politics, bilateral trade, insurance, and information acquisition.
keywords:
Decision-making Under Uncertainty; Risk Aversion; Robust Predictions; Comparative Statics
JEL Classifications: D0; D8
1 Introduction
There are many situations in which one agent must decide which option to present to another in the presence of uncertainty. Consider, for instance, a consumer deciding between firm $1$ and firm $2$’s products, without knowing which best suits her needs. Firm $2$ produces product $b$, while firm $1$ can choose to produce either product $a$ or product $\hat{a}$. If the consumer prefers $a$ to $b$, what are the properties of $\hat{a}$ that guarantee that the consumer also prefers it to $b$? What if firm $1$ knows neither the consumer’s subjective belief nor her utility function?
To elaborate, suppose firm $1$ and firm $2$ are auto manufacturers. Each firm’s car performs differently in different driving conditions, offering one monetary payoff to the consumer when driven in the city and another when driven on the highway. Firm $1$ knows that the consumer prefers product $a$ to product $b$, but knows neither the consumer’s belief about her likelihood of city versus highway driving nor her utility function. What properties must an alternative product $\hat{a}$ have for the consumer to also prefer it to $b$, given that she prefers $a$?
The same exercise can be conducted in the context of political elections, where party $A$ must choose between candidates $a$ and $\hat{a}$ to run against party $B$’s candidate $b$. When would party $A$ benefit more from running candidate $\hat{a}$ over candidate $a$, given that $b$ is the opponent? Allowing for a population with wildly heterogeneous beliefs and utility functions, what are the characteristics of $\hat{a}$ that would ensure at least as many people choose it over $b$ as they would $a$?
The knee-jerk response to these questions is that it is obvious: when the decision-maker’s (DM’s)–i.e., consumer’s or voter’s–utility is known only to be within the class of increasing-in-money utility functions, it must be that $\hat{a}$ corresponds to a first-order stochastic dominance improvement over $a$. Likewise, when the DM is known also to be risk averse, it must be that $\hat{a}$ corresponds to a second-order stochastic dominance improvement over $a$. However, these answers do not stand up to scrutiny–in particular, we do not know the lotteries produced by the DM’s subjective belief, so we cannot speak directly of dominance of lotteries. [^1] In fact, as we reveal in §$\S$4.1, even if we know the the DM’s belief, a dominance improvement of the lottery induced by $a$ is sufficient, but not necessary–a weaker condition suffices. State-wise dominance of $a$ or $b$ is also a promising answer, but it turns out that that is too strong when the DM is known to be risk averse.
In our two main results, Theorems 3.1 and 3.6, we fully characterize these relations in terms of the primitives of the environment–the state-dependent payoffs to actions $a$, $\hat{a}$, and $b$. Theorem 3.1 concerns the case of a risk-averse DM. We show that the DM preferring $a$ to $b$ must imply that she also prefers $\hat{a}$ to $b$ if and only if $\hat{a}$ dominates an action whose payoffs are a convex combination of the payoffs to $a$ and $b$. If $a$ is preferred to $b$, then mixing some of $a$ with $b$ is better than $b$ as well. Then, making $\hat{a}$ dominate such a mixture can only further increase its attractiveness versus $b$.
If we do not assume the DM to be risk averse, our other main result (Theorem 3.6) states that the necessary and sufficient conditions become much more restrictive: $\hat{a}$ dominates $a$ or $b$ (in a state-wise sense). If $u$ is concave, we know from Theorem 3.1 that significant structure on payoff transformations is required. Convexity also imposes structure, but in the opposite manner. Combining these requirements leaves only dominance as necessary (and it is obviously sufficient).
With these results in hand, we discuss several examples. Up first is a political example: in a two-party election, suppose a party is contemplating whether to replace its nominee, candidate $a$. Given that the opponent is candidate $b$, what are the properties of a candidate that is robustly more appealing versus $b$ than $a$? Our first theorem applies directly, so that she must dominate a convex combination of $a$ and $b$ in the sense of our first theorem. Our other two examples posit a status quo trading arrangement–a trade of a risky asset, or a sale of an insurance contract–and asks what contracts must be accepted given that the status quo contract is accepted. Again, our main result tells us that they must be convex combinations of sorts (now, with respect to the outside options).
We finish the paper by studying four natural variants of our main question. Theorem 3.1 assumes two dimensions of uncertainty about the protagonist; both her belief and utility function. What if we shut each of the two down, in turn? In §$\S$4.1, we posit that the DM’s belief is known then characterize, in Theorem 4.2, “how to make a lottery better.” Symmetrically, §$\S$4.2 covers the case with known utilities but unknown beliefs. §$\S$4.3 restores both dimensions of uncertainty; there, we ask for robust improvements to an action vis-a-vis not just one but multiple alternatives. In Proposition 4.7, we show that in a broad class of decision problems, a straightforward aggregation of Theorem 3.1’s conditions are necessary and sufficient for one action to improve against many.
Our final section endogenizes the DM’s belief. Namely, §$\S$5 explores information acquisition, in which our question becomes: how must $a$ relate to $\hat{a}$ in a way that guarantees that $\hat{a}$ is chosen more frequently in the DM’s problem of binary-choice with flexible information acquisition? With two states, we show that $\hat{a}$ dominating $a$ or $b$ is necessary and sufficient. Notably, we illustrate that with three or more states, dominating $a$ is no longer sufficient. That is, $\hat{a}$ can be made better than $a$ in a strong sense (state-wise dominance), yet be chosen less.
1.1 Related Work
The body of work studying decision-making under uncertainty is sizeable. The work closest to this one is Pease and Whitmeyer (2023). There, we formulate a binary relation between actions: action $a$ is safer than $b$ if the the set of beliefs at which $a$ is preferred to $b$ grows larger, in a set inclusion sense, when we make the DM more risk averse. Both Whitmeyer (2024) and Whitmeyer (2025) built off of this paper: the former revisits Rabin (2000)’s calibration exercise in a subjective world, and the latter studies robust comparative statics.
Rothschild and Stiglitz (1970) is a seminal work that characterizes (mean-preserving) transformations of lotteries that are preferred by all risk-averse agents. Aumann and Serrano (2008) formulate an “measure of riskiness” of gambles, as do Foster and Hart (2009) (who are subsequently followed up upon by Bali, Cakici, and Chabi-Yo (2011) and Riedel and Hellmann (2015)). Crucially, these indices and measures correspond to inherently stochastic objects–the lotteries at hand. Our conception of an improvement to an action concerns comparisons of state-dependent payoffs, which are themselves non-random objects (they are just real numbers).
Naturally this work is also connected to the broader literature studying actions that are comparatively friendly toward risk. In addition to the aforementioned paper of ours, Pease and Whitmeyer (2023), which, like this one, centers around a decision-maker’s state-dependent payoffs to actions, this literature includes Hammond III (1974), Lambert and Hey (1979), Karlin and Novikoff (1963), Jewitt (1987), and Jewitt (1989). Notably, they are statements about lotteries, viz., random objects.
Our research question can be reformulated as a revealed-preference exercise. This equivalence is central to deriving our result in the known-beliefs portion of the paper (§$\S$4.2), where we use a remarkable result of Fishburn (1975) “off the shelf.” Existing results are not as useful in our main environment–with uncertain beliefs and utilities. Nevertheless, a number of papers bear mention. Gilboa and Samuelson (2022) ask when beliefs can justify a collection of prize and certainty-equivalent pairs. [^2] Their main result is similar to the result we use from Fishburn (1975). This is due to the fact that both papers’ questions boil down to whether particular separating hyperplanes exist, which is (of course) a common theme running through the revealed-preference literature. Richter and Shapiro (1978) scrutinize what a collection of binary choices between actions reveal about an agent’s beliefs. Echenique and Saito (2015) characterize the market behavior of a risk-averse subjective expected utility maximizing agent, showing that consistent behavior is equivalent to satisfying the “strong axiom of revealed subjective utility.”
Ours is vaguely a comparative statics work–we’re changing an aspect of a decision problem and seeing how it affects a decision-maker’s choice. Our robustness criterion as well as the simplicity of our setting distinguishes our work from the standard pieces, e.g., Milgrom and Shannon (1994), Edlin and Shannon (1998), and Athey (2002). The works involving aggregation (Quah and Strulovici (2012), Choi and Smith (2017), and Kartik, Lee, and Rappoport (2023)) are closer still–as this inherently corresponds to distributional robustness–but none leave as free parameters both the distribution over states and the DM’s utility function, as we do. Special mention is due to Curello and Sinander (2019), who conduct a robust comparative statics exercise in which an analyst, with limited knowledge of an agent’s preferences, predicts the agent’s choice across menus.
In our final section, we explore properties of the new action, $\hat{a}$, that make it more likely to be selected than $a$ when the alternative is $b$ if information is endogenously acquired. This property is similar in spirit to the observation of Matějka and McKay (2015) that new actions added to a menu may “activate” previously un-chosen actions. Muller-Itten, Armenter, and Stangebye (2021) provide a full characterization of this phenomenon. One crucial distinction between our analysis and theirs is that they explore additions whereas our modification is a replacement.
2 Model
There is a topological space of states, $\Theta$, which is endowed with the Borel $\sigma$-algebra, and which we assume to be compact and metrizable. $\theta$ denotes a generic element of $\Theta$. We denote the set of all Borel probability measures on $\Theta$ by $\Delta\equiv\Delta\left(\Theta\right)$. There is also a decision-maker (DM), who is endowed with two actions, $a$ and $b$. $A=\left\{a,b\right\}$ denotes the set of actions, and each action $\tilde{a}\in A$ is a continuous function from the state space to the set of outcomes, $\tilde{a}\colon\Theta\to\mathbb{R}$. For convenience, for any $\tilde{a}\in A$, we write $\tilde{a}_{\theta}\equiv\tilde{a}\left(\theta\right)$. Given a probability distribution over states $\mu\in\Delta$, an action is a (simple) lottery.
We further specify that no action $\tilde{a}\in A$ is weakly dominated by the other. Therefore, we can partition $\Theta$ into three sets:
$\begin{split}\mathcal{A}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>b_{\theta}\right\},\quad\mathcal{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=b_{\theta}\right\}\text{.}\end{split}$
The DM is an expected-utility maximizer, with a von Neumann-Morgenstern utility function defined on the outcome space $u\colon\mathbb{R}\to\mathbb{R}$. We posit that $u$ is strictly increasing, weakly concave, and continuous. On occasion, we will drop the assumption that $u$ is weakly concave, merely requiring that it be strictly increasing (and continuous).
Given $A$, we are interested in how a modification of $a$ affects the DM’s choice of action. That is, we will modify $a$ to some new $\hat{a}$, in which case the new menu is $\left(\hat{a},b\right)$, and examine when $\hat{a}$ must be (strictly) preferred to $b$ whenever $a$ is. Formally,
Definition 2.1.
Action $\hat{a}$ is $b$-Superior to action $a$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
[…]
3 Robust Improvements
We now present our main result, characterizing the conditions under which $\hat{a}$ is b-Superior to $a$. To do so, we first define a Mixture of actions $a$ and $b$ to be the action $a^{\lambda}$ that yields payoff
$a^{\lambda}_{\theta}\coloneqq\lambda a_{\theta}+\left(1-\lambda\right)b_{\theta}$
in each state $\theta\in\Theta$ for some $\lambda\in\left[0,1\right]$.
Our main result is, understanding dominance in a state-wise sense,
Theorem 3.1.
Fix $a$ and $b$. Action $\hat{a}$ is $b$-superior to action $a$ if and only if $\hat{a}_{\theta}>b_{\theta}$ for all $\theta\in\mathcal{A}$ and $\hat{a}$ (weakly) dominates a mixture of $a$ and $b$.
Proof.
[…]
By independence (which is implied by expected utility), for any $\lambda\in\left[0,1\right]$,
$L^{\lambda}\coloneqq\lambda L_{a}+\left(1-\lambda\right)L_{b}\succeq L_{b}\text{.}$
Moreover, by assumption there is some $\lambda^{*}\in\left[0,1\right]$ for which $\lambda^{*}a_{\theta}+\left(1-\lambda^{*}\right)b_{\theta}=\hat{a}_{\theta}$ for all $\theta\in\Theta$. This implies that $L^{\lambda^{*}}$ is a mean preserving spread of $L_{\hat{a}}$. Consequently, as the DM is risk averse, $L_{\hat{a}}\succeq L^{\lambda^{*}}\succeq L_{b}$.∎
The necessity proof of Theorem 3.1 is a little more involved. We prove it in two steps. [^4] Mimicking the approach taken in Jewitt (1986) produces a shorter direct proof of necessity–and we thank a perceptive reviewer for alerting us to this. However, these lemmas will prove useful later on (in our proof of Proposition 4.7), so it is worth temporarily suffering a slight inefficiency. We say that $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$ if for any pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$, there exists a $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta^{\prime}}\text{.}$
The first of the two necessity steps is showing that $b$-superiority implies pairwise dominance.
Lemma 3.2.
$\hat{a}$ is $b$-superior to $a$ only if i) for any $\theta^{\dagger}\in\mathcal{C}$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$; ii) for any $\theta\in\mathcal{A}$, $\hat{a}_{\theta}>b_{\theta}$; and iii) $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$.
Proof.
See Appendix A. ∎
Let us consider the meaning of Lemma 3.2. First, it is clear that if $\hat{a}$ is $b$- superior, then i) and ii) must be true. Next, if iii) is not true and if $\hat{a}_{\theta^{\prime}}<a_{\theta^{\prime}}<b_{\theta^{\prime}}$ and $\hat{a}_{\theta}>a_{\theta}>b_{\theta}$, there are beliefs–assigning zero probability to all states other than the specified $\theta$ and $\theta^{\prime}$–at which a risk-neutral DM will rank the actions $a\succ b\succ\hat{a}$, so $\hat{a}$ is not $b$-superior. On the other hand, if these two chains do hold, although $a\succeq b$ will always imply $\hat{a}\succeq b$ for a risk-neutral DM, we can nevertheless always construct a (concave) utility function such that there exist–again, with positive probability assigned only to states $\theta$ and $\theta^{\prime}$–beliefs that produce $a\succ b\succ\hat{a}$.
Finally, we show that pairwise domination implies domination of a mixture, closing the proof of necessity.
Lemma 3.3.
If $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$ for all $\theta^{\dagger}\in\mathcal{C}$, and $\hat{a}_{\theta}\geq b_{\theta}$ for all $\theta\in\mathcal{A}$, then $\hat{a}$ dominates a mixture of $a$ and $b$.
Proof.
Suppose $\hat{a}$ pairwise-dominates a collection of mixtures of $a$ and $b$, $\hat{a}_{\theta^{\dagger}}\geq a_{\theta^{\dagger}}$ for all $\theta^{\dagger}\in\mathcal{C}$, and $\hat{a}_{\theta}\geq b_{\theta}$ for all $\theta\in\mathcal{A}$. This implies that for any pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$,
$\min\left\{1,\frac{\hat{a}_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}}\right\}\geq\lambda_{\theta,\theta^{\prime}}\geq\max\left\{0,\frac{b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}{b_{\theta^{\prime}}-a_{\theta^{\prime}}}\right\}\text{.}$
This implies that there exists a $\lambda\in\left[0,1\right]$ for which
$\inf_{\theta\in\mathcal{A}}\left\{1,\frac{\hat{a}_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}}\right\}\geq\lambda\geq\sup_{\theta^{\prime}\in\mathcal{B}}\left\{0,\frac{b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}{b_{\theta^{\prime}}-a_{\theta^{\prime}}}\right\}\text{,}$
and so $\hat{a}$ dominates a mixture of $a$ and $b$.∎
We note that Theorem 3.1 can easily be extended to the case in which a lower-bound (in terms of risk-aversion) for the DM is known. That is, suppose there is some (weakly) risk-averse $\bar{u}$ of which the DM’s utility is known to be some strictly increasing (weakly) concave transformation: $u=\phi\circ\bar{u}$. Then, we just redefine $a$ from map $\theta\mapsto a(\theta)$ to $\theta\mapsto\bar{u}(a(\theta))$ and likewise $b$, before applying the theorem. Similarly, as second-order stochastic dominance is preserved under independent noise, [^5] See, for instance, the discussion on page 3 of Pomatto, Strack, and Tamuz (2020). the theorem also holds when we ask for robustness to aggregate risk.
3.1 Intuition
To gain intuition for Theorem 3.1, consider the two-state environment, which is easy to visualize. The state is either 00 or $1$, and we specify without loss of generality that $a_{0}>b_{0}$ and $b_{1}>a_{1}$. $\mu\in\left[0,1\right]$ denotes the DM’s belief that the state is $1$.
There is a cutoff belief at which the DM is indifferent between actions. Let $\bar{\mu}_{u}$ be the belief at which the DM is indifferent between actions $a$ and $b$ with utility $u$ and $\hat{\mu}_{u}$ be the belief at which the DM is indifferent between $\hat{a}$ and $b$ with the same utility. To see this graphically, let
$\ell_{a}(\mu)=a_{0}(1-\mu)+a_{1}\mu\text{,}$
be the expected payoff (in money) of action $a$ as a function of the belief $\mu\in\left[0,1\right]$. In turn, lines
$\ell_{\hat{a}}(\mu)=\hat{a}_{0}(1-\mu)+\hat{a}_{1}\mu\quad\text{and}\quad\ell_{b}(\mu)=b_{0}(1-\mu)+b\mu\text{,}$
are the expected payoffs to actions $\hat{a}$ and $b$, respectively. In Figure 1, these are the blue (solid), red (dashed), and green (dotted) lines. The top panel of the figure shows the monetary payoffs, or the risk-neutral utility function, while the bottom panel depicts utilities with $u\left(\cdot\right)=\sqrt{1+\cdot}-3$. Then, $\hat{a}$ being $b$-superior to $a$ is equivalent to $\bar{\mu}_{u}\leq\hat{\mu}_{u}$ for all concave $u$.
Figure 1: Improving $a$. Try it yourself (rotate by moving the slider)!
Now we can think about $b$-superiority in the context of Theorem 3.1. It tells us that $\hat{a}$ must at least weakly dominate a mixture of $a$ and $b$. For now, just consider an $\hat{a}$ that is a mixture of $a$ and $b$. The corresponding line, $\ell_{\hat{a}}$, is a counterclockwise rotation of $\ell_{a}$. We know from the theorem that for any belief at which the DM preferred $a$ to $b$ ($\mu\leq\bar{\mu}_{u}$), she will prefer $\hat{a}$ to $b$ as well. In other words, if $a$ was better than $b$ at $\mu$, then mixing “some” $a$ with $b$ is better than $b$ at $\mu$ as well. Consequently, the set of beliefs at which $\hat{a}$ is preferred must be at least as large as the set where $a$ was preferred. Naturally, making $\hat{a}$ better than a mixture of $a$ and $b$ only strengthens its attractiveness versus $b$.
Formally, we can write the following result.
Corollary 3.4.
$\hat{a}$ is $b$-superior to $a$ if and only if $\hat{a}_{0}>b_{0}$ and there exists a $\lambda\in\left[0,1\right]$ such that $\ell_{\hat{a}}$ lies weakly above the line $\lambda\ell_{a}+\left(1-\lambda\right)\ell_{b}$.
3.2 The Toothlessness of Expected Utility
Before going on, we wish to highlight one surprising feature of our theorem; namely, the fact that the sufficiency direction–and, therefore, the entire Theorem 3.1–is valid for far more general preferences than expected utility. We understand an action as a vector in $\mathbb{R}^{\Theta}$ and suppose that a DM’s preferences are represented by the monotone and concave $V\colon\mathbb{R}^{\Theta}\to\mathbb{R}$, where she (strictly) prefers $a$ to $b$ if and only if $V(a)\geq\ (>)\ V(b)$.
Suppose that $\hat{a}$ dominates a mixture of $a$ and $b$, $\hat{a}\neq b$, and the DM (strictly) prefers $a$ to $b$. Then, for some $\lambda\in\left(0,1\right]$, by concavity and monotonicity,
$V\left(\hat{a}\right)\geq V\left(\lambda a+(1-\lambda)b\right)\geq\lambda V(a)+(1-\lambda)V(b)\underset{(>)}{\geq}V(b)\text{.}$
In fact, if we dropped the requirement that strict preferences are preserved, we could weaken the requirement on $V$ to be merely quasi-concave.
3.3 A Broader Class of Utilities
In the work up to this point, we have assumed mild but meaningful structure on the DM’s utility function. More is better ($u$ is strictly increasing in the reward), and the DM is risk-averse ($u$ is, at least weakly, concave). Suppose we remove the assumption of risk aversion. We still assume that $u$ is strictly increasing but now allow it to take arbitrary shape.
We then define the equivalent notion to $b$-superiority without risk aversion.
Definition 3.5.
Action $\hat{a}$ is $b$-Better than action $a$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, but not necessarily concave, continuous $u$.
As we now document, without the assumption of risk aversion, the only transformations that must make $b$ less attractive are unambiguous dominance improvements from $a$ to $\hat{a}$. This is intuitive; think back to Corollary 3.4. There, the necessary relationship between $\hat{a}$ and $a$ is $\ell_{\hat{a}}$’s dominance of a counterclockwise rotation of $\ell_{a}$. However, the opposite modification is needed when $u$ is convex–which is now permitted–that is, $\ell_{\hat{a}}$ needs to dominate a clockwise rotation of $\ell_{a}$. This leaves only dominance improvements.
Theorem 3.6.
Action $\hat{a}$ is $b$-better than action $a$ if and only if for all $\theta\in\mathcal{A}$ $\hat{a}_{\theta}>b_{\theta}$ and $\hat{a}$ dominates $a$ or $b$.
Proof.
See Appendix A. ∎
3.4 Applications
3.4.1 Choosing Which Candidate to Run
There is a unit mass of voters and two candidates $a$ and $b$ running for office. There is an unknown state of the world $\theta\in\Theta$, and candidate $a$ ($b$) produces monetary benefit $a_{\theta}$ ($b_{\theta}$) to each voter in every state $\theta\in\Theta$. Neither candidate dominates the other in the sense that there is at least one state of the world in which each provides a strictly higher monetary benefit to the populace than the other.
The party backing candidate $a$ is contemplating whether to replace him with a different candidate $\hat{a}$. Crucially, it knows neither the risk preferences of the populace (which may be heterogeneous) nor the beliefs of the populace (also possibly heterogeneous). Accordingly, it makes its choice based on a robust criterion, asking what are the candidates who have a better chance of winning versus $b$, no matter the population’s (concave) utilities or beliefs. Who cannot be worse than $a$ versus $b$? Theorem 3.1 tells us the answer:
Remark 3.7.
Candidate $\hat{a}$ is no worse than $a$ versus $b$ if and only if she dominates a mixture of $a$ and $b$.
In the context of politics, this can be understood, roughly, as a move to the political center. This strategy has been explicitly used in the past. For example, after Republican Ronald Reagan’s landslide presidential victory in 1984, a group of Democrats founded the Democratic Leadership Council (DLC) in 1985 with the explicit goal of shifting the party towards the center in order to win elections. Likewise, Tony Blair’s “New Labour” and the CDU’s adoption of Green Party’s environmental policies are both similar convexifications.
3.4.2 Robustly Optimal Bilateral Trade Modifications
There is a buyer ($B$) and a seller ($S$). $S$ possesses an asset that pays out $v_{\theta}\in\mathbb{R}$ in state $\theta$. The state is contractible and the status quo trade agreement sees transfers of size $\gamma_{\theta}\in\mathbb{R}$ from $B$ to $S$ in each state $\theta\in\Theta$. We assume that there exist states $\theta^{\prime},\theta^{\prime\prime}\in\Theta$ for which $v_{\theta^{\prime}}>\gamma_{\theta^{\prime}}$ and $\gamma_{\theta^{\prime\prime}}>v_{\theta^{\prime\prime}}$.
$B$’s state-dependent payoff from transacting is $v_{\theta}-\gamma_{\theta}$, and $S$’s is $\gamma_{\theta}$ (budget balance). $B$’s outside option is the sure-thing 00 and $S$’s outside option is the asset, with random payoff $v_{\theta}$ in each state $\theta\in\Theta$.
We suppose that the status-quo is acceptable–given the agents’ subjective beliefs about uncertainty and their (concave) utility functions, each is willing to participate in the arrangement. What modifications to the arrangement are robustly optimal in that $B$ and $S$ will still remain willing to participate? Theorem 3.1 reveals the answer.
We observe that for all $\theta\in\Theta$ and for any $\lambda\in\left[0,1\right]$,
$\lambda\left(v_{\theta}-\gamma_{\theta}\right)=v_{\theta}-\lambda\gamma_{\theta}-\left(1-\lambda\right)v_{\theta}\text{,}$
so that any new transfer that yields the requisite convex combination of $B$’s status quo payoff from accepting the terms and $B$’s outside option of 00 equals
$\lambda\gamma_{\theta}+\left(1-\lambda\right)v_{\theta}\text{,}$
which is precisely the convex combination (with the same weight) of $S$’s status quo payoff and her outside option. This, plus our assumption of budget balance, imply
Remark 3.8.
A new trade agreement, $\left(\hat{\gamma}_{\theta}\right)_{\theta\in\Theta}$ must be acceptable if and only if there exists some $\lambda\in\left(0,1\right]$ such that $\hat{\gamma}_{\theta}=\lambda\gamma_{\theta}+\left(1-\lambda\right)v_{\theta}$ for all $\theta\in\Theta$.
3.4.3 Robustly Optimal Insurance Modifications
Now consider the scenario of a risk-neutral insurer and a consumer. The consumer’s payoff without insurance is $v_{\theta}$ in each state $\theta\in\Theta$. The status-quo policy yields a payoff (net of the loss) of $\alpha_{\theta}$ to the consumer in each state $\theta\in\Theta$. We assume that there exist some $\theta^{\prime},\theta^{\prime\prime}\in\Theta$ for which $\alpha_{\theta^{\prime}}>v_{\theta^{\prime}}$ and $\alpha_{\theta^{\prime\prime}}<v_{\theta^{\prime\prime}}$.
Remark 3.9.
A new contract, $\left(\hat{\alpha}_{\theta}\right)_{\theta\in\Theta}$ must be acceptable if and only if there exists some $\lambda\in\left(0,1\right]$ such that $\hat{\alpha}_{\theta}\geq\lambda\alpha_{\theta}+\left(1-\lambda\right)v_{\theta}$ for all $\theta\in\Theta$.
One interesting implication of this remark is that the set of acceptable contracts could very well expose the consumer to more risk. Indeed, suppose that the status-quo policy is risk free, i.e., yields the consumer the same wealth in every state. The consumer’s outside option (not buying insurance) is, of course, riskier. In this case, any actuarily-fair contract that the consumer must accept also exposes her to risk as her wealth is state-dependent.
4 Extensions and Generalizations
In this section, we look beyond pairwise comparisons where we are uncertain about both the DM’s belief and her utility. In particular, we study how to make an action better versus multiple alternatives, not just one, in three different environments. First, we assume that the DM’s belief is known, shutting down one of the two dimensions of uncertainty. What makes a lottery better? Second, we suppose that we know the DM’s utility but not her belief. Third, we restore the main specification in which we seek robustness with respect to both beliefs and utilities.
4.1 How to Make a Lottery Better
We begin by characterizing how to make an action better when we know a DM’s belief but do not know her utility.
We assume that the DM has a finite set of actions $A$ (with $\left|A\right|=m+1\geq 2$). For any fixed $\mu$, each action $a\in A$ induces a lottery $L_{a}$. For any $a\in A$, set $B=A\setminus\left\{a\right\}$ denotes the DM’s set of actions other than $a$. We also let $\succ$ denote the DM’s preference relation over lotteries that corresponds to her utility function. Our definition of a robust improvement in the known-lottery case is, therefore,
Definition 4.1.
Lottery $L_{\hat{a}}$ $B$-Improves upon lottery $L_{a}$ if
$L_{a}\succ L_{b}\text{ for all }b\in B\ \Rightarrow\ L_{\hat{a}}\succeq L_{b}\text{ for all }b\in B\text{,}$
for all utilities in the specified class.
Concordant with the rest of this paper, our two leading classes of utility functions are those for a risk-averse DM–continuous, strictly increasing, and concave–and a monotone DM–continuous, and strictly increasing. The details of this definition are slightly inconsistent with the other definitions in this paper; e.g., $a$ is strictly preferred to the other actions, whereas we ask only that $\hat{a}$ be weakly preferred to the others. We do this only for simplicity: this formulation allows for an especially clean result. [^6] To elaborate, our approach makes use of the revealed-preference results of Fishburn (1975). Our definition allows us to use his Corollary 3, in which each revealed preference is strict. Otherwise, we could appeal to Fishburn (1975)’s result for the “finite composite case,” Theorem 3, which would produce a messier result than what we have (Theorem 4.2).
[…]
Theorem 4.2.
Lottery $L_{\hat{a}}$ $B$-improves upon lottery $L_{a}$ if and only if for all $b\in\left\{1,\dots,m\right\}$ there exists a convex weight $\lambda$ such that
$\sum_{i=1}^{m}\lambda_{i}L_{i}+\left(1-\sum_{i=1}^{m}\lambda_{i}\right)L_{\hat{a}}\trianglerighteq\sum_{i=1}^{m}\lambda_{i}L_{a}+\left(1-\sum_{i=1}^{m}\lambda_{i}\right)L_{b}\text{.}$
Proof.
We observe that lottery $L_{\hat{a}}$ $B$-improving upon lottery $L_{a}$ is equivalent to the statement that there does not exist a utility function in the specified class such that
$L_{a}\succ L_{b}\text{ for all }b\in B,\text{ and }L_{b}\succ L_{\hat{a}}\text{ for some }b\in B\text{.}$
By Corollary 3 in Fishburn (1975), this holds if and only if for all $b\in B$
$\sum_{i=1}^{m}\lambda_{i}L_{i}+\left(1-\sum_{i=1}^{m}\lambda_{i}\right)L_{\hat{a}}\trianglerighteq\sum_{i=1}^{m}\lambda_{i}L_{a}+\left(1-\sum_{i=1}^{m}\lambda_{i}\right)L_{b}\text{,}$
for some convex weight $\lambda$.∎
When there are just two actions (so $B=\left\{b\right\}$), we alter the definition of improvement in the obvious way and have
Corollary 4.3.
Lottery $L_{\hat{a}}$ $b$-improves upon lottery $L_{a}$ if and only if there exists $\lambda\in\left[0,1\right]$ such that
$\lambda L_{b}+\left(1-\lambda\right)L_{\hat{a}}\trianglerighteq\lambda L_{a}+\left(1-\lambda\right)L_{b}\text{.}$
Fixing the DM’s belief in the revealed-preference exercise results in a linear problem, which allows us to use tools that we are unable to employ otherwise. Specifically, Corollary 3 in Fishburn (1975)–and therefore Theorem 4.2–is implied by a separating hyperplane theorem. This produces our observation that if $\hat{a}$ does not $B$-improve on $a$, then we cannot find a combination of $\hat{a}$ and $b$ that dominates a combination of $b$ and $a$.
4.2 Known Utilities
Now suppose, symmetric to §$\S$4.1, that we do not know the DM’s belief, but instead know her utility function. Thus, we assume without loss of generality that $u\left(\cdot\right)=\cdot$, i.e., that the DM is risk-neutral. We maintain the specification that $A$ is finite, containing at least two actions; and to avoid trivialities, we stipulate further that no action in $A$ is weakly dominated.
As there are $n\geq 2$ states, $\Delta$ is the $\left(n-1\right)$-dimensional probability simplex. Moreover, our assumptions mean that the region of beliefs on which any specified action $a\in A$ is optimal is an $\left(n-1\right)$-dimensional polytope $P_{a}$:
$P_{a}\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}a_{\theta}\geq\max_{b\in A}\mathbb{E}_{\mu}b_{\theta}\right\}\text{.}$
We also define the set of beliefs where the DM is indifferent between $a$ and $b$ for each $b\in A\setminus\left\{a\right\}$, or
$H_{b}^{a}\coloneqq\left\{\mu\in\Delta\colon\mathbb{E}_{\mu}a_{\theta}=\mathbb{E}_{\mu}b_{\theta}\right\}\text{,}$
and say that an action $b$ is $a$-Relevant if
$\dim\left\{H_{b}^{a}\cap P_{a}\right\}=n-2\text{.}$
For any $a$-relevant $b$, we let $F_{b}^{a}\coloneqq H_{b}^{a}\cap P_{a}$. $B_{a}\subseteq A$ denotes the set of $a$-relevant actions.
We let $V_{a}$ denote the set of vertices (00-faces) of $P_{a}$ and let $V_{b}^{a}$ denote the set of vertices (00-faces) of $F_{b}^{a}$. We say that $\hat{a}$ is a $b$-Rotation around a vertex if for all $\mu\in V_{b}^{a}$, $\mathbb{E}_{\mu}\hat{a}_{\theta}\geq\mathbb{E}_{\mu}a_{\theta}$. We also observe that $\cup_{b\in B_{a}}V_{b}^{a}\subseteq V_{a}$, with equality if and only if there does not exist a state in which $a$ is uniquely optimal.
Definition 4.4.
Action $\hat{a}$ $u$-improves upon action $a$ if
$\mathbb{E}_{\mu}a_{\theta}\geq\max_{b\in A}\mathbb{E}_{\mu}b_{\theta}\quad\Rightarrow\quad\mathbb{E}_{\mu}\hat{a}_{\theta}\geq\max_{b\in A}\mathbb{E}_{\mu}b_{\theta}\text{,}$
and
$\mathbb{E}_{\mu}a_{\theta}>\max_{b\in B}\mathbb{E}_{\mu}b_{\theta}\quad\Rightarrow\quad\mathbb{E}_{\mu}\hat{a}_{\theta}>\max_{b\in B}\mathbb{E}_{\mu}b_{\theta}\text{.}$
Then, recalling that $\mathcal{A}$ is the set of states in which $a$ is uniquely optimal,
Theorem 4.5.
Action $\hat{a}$ $u$-improves upon action $a$ if and only if for all $\theta\in\mathcal{A}$, $\hat{a}_{\theta}>\max_{b\in B}b_{\theta}$, and for any $a$-relevant $b$, $\hat{a}$ is a $b$-rotation around a vertex.
Proof.
Necessity is immediate. As for sufficiency, observe that $P_{a}$ is a compact convex subset of $\Delta$, so by the the Krein-Milman theorem, it is the closed convex hull (cch) of its extreme points. Specifically, $P_{a}=\operatorname{cch}V_{a}$. By construction, for each $\mu\in V_{a}$, $\mathbb{E}_{\mu}\hat{a}_{\theta}\geq\mathbb{E}_{\mu}a_{\theta}$, so $\mathbb{E}_{\mu}\hat{a}_{\theta}\geq\mathbb{E}_{\mu}a_{\theta}\geq\max_{b\in A}\mathbb{E}_{\mu}b_{\theta}$ for all $\mu\in\operatorname{cch}V_{a}=P_{a}$. Finally, by the linearity in $\mu$ of the expectation, for all $\mu\in P_{a}$ such that $\mathbb{E}_{\mu}a_{\theta}>\max_{b\in B}\mathbb{E}_{\mu}b_{\theta}$, $\mathbb{E}_{\mu}\hat{a}_{\theta}>\max_{b\in B}\mathbb{E}_{\mu}b_{\theta}$. ∎
4.3 One Versus Many
We finish by restoring both dimensions of uncertainty. We look for improvements to $a$ that make it more attractive in comparison to multiple alternatives, irrespective of the DM’s belief and utility. Consider our leading example, in which the manager of firm $1$ is deciding between products $a$ and $\hat{a}$, but now the firm is competing against a product from firm $2$ as well as a product from firm $3$. How does the inclusion of firm $3$ affect the manager’s decision?
The main difficulty with considering multiple alternatives in general is that for different levels of risk aversion, the DM’s “next-best” action may be different. To illustrate, consider the following example with actions $a$, $b$, and $c$ and two states $\theta\in\{0,1\}$. Let $c_{0}=c_{1}=2/5$, $a_{1}=b_{0}=0$, and $a_{0}=b_{1}=1$. Then if the DM is risk-neutral, only $a$ and $b$ will ever be chosen for any belief. If, however, the DM is more risk-averse, say $u(x)=\sqrt{x}$, then for the most uncertain beliefs, $c$ is the best choice. This complicates the analysis of the multi-alternative environment significantly because the set of relevant alternatives depends on the level of risk aversion.
Two additional assumptions make our analysis of multiple alternatives tractable. We continue to assume that $A$ is finite, with $m+1\geq 2$ actions and maintain our convention that for a given $a$, $B=A\setminus\left\{a\right\}$. First, we assume that the state space is Rich: for each action $a\in A$, there exists at least one state $\theta\in\Theta$ such that $a_{\theta}>\max_{b\in B}b_{\theta}$–that is, each action is uniquely optimal in at least one state. Second, we assume that preferences take the following Single-Peaked form: for each action $a\in A$, and any pair of states $\theta,\theta^{\prime}\in\Theta$, either
  1. 1.
    there exists $\lambda\in\left[0,1\right]$ such that
    $\lambda a_{\theta}+(1-\lambda)a_{\theta^{\prime}}\geq\max_{b\in B}\left\{\lambda b_{\theta}+(1-\lambda)b_{\theta^{\prime}}\right\},\quad\text{or}$
  2. 2.
    there exists $b\in B$ such that $b_{\theta}\geq a_{\theta}$ and $b_{\theta^{\prime}}\geq a_{\theta^{\prime}}$.
The richness specification is straightforward. The single-peaked requirement is more subtle. For any action $a$ and any pair of states, if there exists a belief supported only on those two states that rationalizes that action, then $a$ must always be included in the set of relevant actions. If instead there exists an action that dominates $a$ on those two states, then action $a$ need never be included as a relevant action. Either way, this results in a constant set of relevant actions, regardless of the level of risk aversion.
This set of specifications is satisfied by the oft-encountered “quadratic loss” monetary reward: $\Theta=\left[0,1\right]$, $A\subset\left[0,1\right]$ and $a_{\theta}=-\left(a-\theta\right)^{2}$. It also holds in our earlier setting of just two actions.
Our superiority definition is
Definition 4.6.
Action $\hat{a}$ is $B$-Superior to action $a$ if
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)\geq\max_{b\in B}\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)\geq\max_{b\in B}\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
and
$\mathbb{E}_{\mu}u\left(a_{\theta}\right)>\max_{b\in B}\mathbb{E}_{\mu}u\left(b_{\theta}\right)\ \Rightarrow\ \mathbb{E}_{\mu}u\left(\hat{a}_{\theta}\right)>\max_{b\in B}\mathbb{E}_{\mu}u\left(b_{\theta}\right)\text{,}$
for any strictly increasing, concave, and continuous $u$.
We extend our notation in an intuitive way:
$\begin{split}\mathcal{A}_{B}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>\max_{b\in B}b_{\theta}\right\},\quad\mathcal{B}_{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<\max_{b\in B}b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}_{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=\max_{b\in B}b_{\theta}\right\}\text{.}\end{split}$
Our result is a natural generalization of Theorem 3.1.
Proposition 4.7.
Fix $a$ and $B$. Action $\hat{a}$ is $B$-superior to action $a$ if and only if i) $\hat{a}_{\theta}>\max_{b\in B}b_{\theta}$ for all $\theta\in\mathcal{A}_{B}$; and ii) $\hat{a}$ (weakly) dominates a mixture of $a$ and $b$, for all $b\in B$.
Proof.
Please see Appendix A.3.∎
As we note in the proof, that aggregating Theorem 3.1’s pairwise conditions implies that $\hat{a}$ improves upon $a$ versus a set of actions is relatively simple. That this feature is necessary is more surprising, and this necessity is engendered by the additional richness and single-peakedness assumptions we place on the decision problem.
5 Information Acquisition
Up until this point we have focused on a simple decision problem in which a DM has a belief over the states of the world and chooses an optimal action based on this belief. It is easy to see that $b$-superiority extends to the scenario in which a DM obtains exogenous information before making a decision. That is, $\hat{a}$ is $b$-superior to $a$ if and only if for any $\mu\in\operatorname{int}\Delta$ and any Bayes-plausible (martingale) distribution over posteriors $F$, the DM chooses $\hat{a}$ with a higher probability than she chooses $a$ (where, in both cases, $b$ is chosen with the complimentary probability).
There has been a recent explosion of interest in problems with endogenous information acquisition. Suppose we want to know what the properties of $\hat{a}$ are that make it chosen more versus $b$ than $a$ is when the DM flexibly acquires information before taking a choice.
We begin with the binary-state environment ($\Theta=\left\{0,1\right\}$). We maintain our assumptions that neither $a$ nor $b$ dominates the other and that $a$ is uniquely optimal in state 00 and $b$ in state $1$. Given a prior $\mu_{0}\in\left(0,1\right)$, and defining an agent’s value function, in belief $\mu$, as
$V(\mu)\coloneqq\max_{\tilde{a}\in\left\{a,b\right\}}\mathbb{E}_{\mu}u(\tilde{a}_{\theta})\text{,}$
the DM’s flexible information acquisition problem is
[…]
Any solution $F^{*}$ ($\hat{F}^{*}$) to Program 5 (5) produces an optimal choice probability of action $a$ ($\hat{a}$) by the DM, which we define to be $p_{D,\mu_{0},u}\equiv p$ ($\hat{p}_{D,\mu_{0},u}\equiv\hat{p}$).
Definition 5.1.
$\hat{a}$ is selected more than $a$ if for any Bayes-plausible and UPS $D$, prior $\mu_{0}\in\operatorname{int}\Delta$, and strictly increasing, concave $u$; for any optimal $p$, there exists an optimal choice probability $\hat{p}\geq p$.
Proposition 5.2.
$\hat{a}$ is selected more than $a$ if and only if $\hat{a}$ dominates $a$ or $b$.
Proof.
A full proof may be found in Appendix A.4. ∎
Most of the work in proving the proposition is in the necessity portion. First consider what happens if $\hat{a}$ is not $b$-superior to $a$. Then, we can find a $u$ such that $\hat{p}<\bar{p}$. Take such a $u$, in which case we can then always find a cost function such that $a$ is selected with probability $1$ when the menu is $\left\{a,b\right\}$ but when the menu is $\left\{\hat{a},b\right\}$, either the DM learns in a nontrivial manner, and hence selects $\hat{a}$ with probability strictly less than $1$, or does not learn but selects $b$ with probability $1$.
To understand why $\hat{a}$ must dominate $a$ or $b$, consider how the DM chooses to learn when there are two states and two actions. The DM chooses an $F^{\ast}$ such that there are two posteriors, $\mu_{L}^{a}$ and $\mu_{H}^{a}\geq\mu_{L}^{a}$ such that she chooses $a$ with probability 1 if $\mu_{0}\leq\mu_{L}^{a}$, $b$ with probability 1 if $\mu_{0}\geq\mu_{H}^{a}$, and $a$ with probability $p\in(0,1)$ if $\mu_{0}\in(\mu_{L}^{a},\mu_{H}^{a})$.
Next, take $\hat{a}$ to be such that $\ell_{\hat{a}}$ is a counter-clockwise rotation from $\ell_{a}$ towards $\ell_{b}$ (recall Figure 1). This is equivalent, however, to an increase in learning cost (if payoffs had remained the same) for a risk-neutral DM. A higher marginal cost of information leads to less learning; i.e., $(\mu_{L}^{\hat{a}},\mu_{H}^{\hat{a}})\subseteq(\mu_{L}^{a},\mu_{H}^{a})$. This means that we can find a prior belief such that $\mu_{0}\in(\mu_{L}^{a},\mu_{H}^{a})$ but $\mu_{0}\geq\mu_{H}^{\hat{a}}$. In other words, action $a$ has a positive probability of being chosen when the menu is $\{a,b\}$, but action $\hat{a}$ will never be chosen when the menu is $\{\hat{a},b\}$. Hence, $b$-superiority is clearly not enough to guarantee that $\hat{a}$ is chosen more than $a$.
5.1 Three or More States
Figure 2: The insufficiency of dominance for three or more states
Alas, when there are three states, $\hat{a}$ dominating $a$ no longer implies that $\hat{a}$ is selected more than $a$, as we now illustrate. Let $\Theta=\left\{0,1,2\right\}$ with $\mu^{2}\coloneqq\mathbb{P}(2)$ and $\mu^{1}\coloneqq\mathbb{P}(1)$. We denote $\alpha_{\theta}\equiv u(a_{\theta})$ for all $\theta$. Let $\hat{\alpha}_{2}=3/2=\alpha_{2}+1/2$, $\alpha_{1}=-1=\hat{\alpha}_{1}$ and $\hat{\alpha}_{0}=\alpha_{0}=0$; and $\beta_{\theta}=0$ for all $\theta\in\Theta$. Evidently, $\hat{a}$ weakly dominates $a$ (and $\hat{a}\neq a$).
We have
$V(\mu^{1},\mu^{2})=\max\left\{\mu^{2}-\mu^{1},0\right\},\quad\text{and}\quad\hat{V}(\mu^{1},\mu^{2})=\max\left\{\frac{3}{2}\mu^{2}-\mu^{1},0\right\}\text{.}$
We then define
$W(\mu^{1},\mu^{2})\coloneqq\max\left\{V,\hat{V}+\frac{1}{4}\mu^{1}-\frac{1}{8}\right\}\text{,}$
and appeal to Lemma A.1 in Whitmeyer (2023), which states that if a set of produced beliefs does not contain multiple posteriors that justify the same action, then there exists a cost function that generates it. We therefore conclude that there exists a UPS cost such that the four points
$\left\{\left(\frac{2}{25},\frac{1}{5}\right),\left(\frac{12}{25},\frac{1}{5}\right),\left(\frac{1}{50},\frac{53}{200}\right),\left(\frac{27}{50},\frac{1}{200}\right)\right\}$
are the support of optimal learning given value function $W$. Accordingly, we take such a cost function and fix prior $\left(3/20,1/5\right)$ then compute $p=33/40>3/4=\hat{p}$. [^8] This calculation simply applies Bayes plausibility to the given posteriors. Specifically, $3/20=2/25p+12/25(1-p)$.
Why is $\hat{a}$ chosen less often, even though it dominates $a$? When the DM faces a choice between $a$ and $b$, she only learns about whether the state is 00 or $1$, as Expression 5.1 reveals. She would like to avoid choosing $a$ in state $1$, so this is the best approach. If, however, the DM has a choice between $\hat{a}$ and $b$, she now learns about all three, either becoming relatively confident that the state is not $1$ ($\hat{a}$ is preferable) or relatively confident that the state is not $2$ ($b$ is preferable). This is enough to make it so that $a$ is chosen more than $\hat{a}$.
Figure 2 illustrates this example on the $2$-simplex. The four colored polygons are the regions in which $V$ lies above the rotated value function $\hat{V}$ and either $a$ is optimal (red) or $b$ is optimal (orange); and the rotated value function $\hat{V}$ lies above $V$ and either $\hat{a}$ is optimal (purple) or $b$ is optimal (blue). The prior is the hollow red dot. The black dots are the support of $F^{*}$, and the orange $x$s are the support of $\hat{F}^{*}$.
Leaving the single-dimensional environment corresponding to the two-state simplex engenders this result. What we are doing is taking a value function that in a sense “twist things,” making it so that the learning in the two problems is not along the same line segment. There is enough freedom then to pick an appropriate prior so that $p>\hat{p}$.
We began this paper with the examples of a firm looking to robustly improve its product versus another, or a political party seeking to enlarge its platform. This example reveals that if beliefs are endogenous, even a seemingly obvious improvement can be harmful. That is, a firm’s product can be made unambiguously better–even via a price decrease–yet the competitor’s product be chosen more. An indisputable upgrade to a political party’s policies can result in a lost election.
Appendix A Omitted Proofs
A.1 Lemma 3.2 Proof
Proof.
The first two conditions are obviously necessary for $b$-superiority. We assume they hold and suppose for the sake of contraposition that $\hat{a}$ does not pairwise-dominate a collection of mixtures of $a$ and $b$. This means that there exists at least one pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}\times\mathcal{B}$ for which there exists no $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that the two inequalities in Expression 3 hold.
Fix such a pair of $\theta$ and $\theta^{\prime}$. Given utility $u$, let $\bar{\mu}_{u}$ solve
$\left(1-\bar{\mu}_{u}\right)u\left(b_{\theta}\right)+\bar{\mu}_{u}u\left(b_{\theta^{\prime}}\right)=\left(1-\bar{\mu}_{u}\right)u\left(a_{\theta}\right)+\bar{\mu}_{u}u\left(a_{\theta^{\prime}}\right)\text{,}$
and $\hat{\mu}_{u}$ solve
$\left(1-\hat{\mu}_{u}\right)u\left(b_{\theta}\right)+\hat{\mu}_{u}u\left(b_{\theta^{\prime}}\right)=\left(1-\hat{\mu}_{u}\right)u\left(\hat{a}_{\theta}\right)+\hat{\mu}_{u}u\left(\hat{a}_{\theta^{\prime}}\right)\text{;}$
viz., these are the beliefs on the edge of $\Delta$ between the the degenerate probability distributions on states $\theta$ and $\theta^{\prime}$, $\delta_{\theta}$ and $\delta_{\theta^{\prime}}$, at which the DM with utility $u$ is indifferent between $a$ and $b$ and $\hat{a}$ and $b$, respectively. If the DM is risk-neutral, we drop the subscript and simply write $\bar{\mu}$ and $\hat{\mu}$.
By assumption $\bar{\mu},\hat{\mu}\in\left(0,1\right)$. Moreover, observe that a necessary condition for $b$-superiority is $\hat{\mu}\geq\bar{\mu}$, which holds if and only if
$\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)\left(\hat{a}_{\theta}-b_{\theta}\right)\geq\left(a_{\theta}-b_{\theta}\right)\left(b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}\right)\text{.}$
If this inequality does not hold, we are done, as we have produced the right negation. Accordingly, suppose it does hold. This means, given our other assumptions,
$b_{\theta}<a_{\theta}<\hat{a}_{\theta},\quad\text{and}\quad\hat{a}_{\theta^{\prime}}<a_{\theta^{\prime}}<b_{\theta^{\prime}}\text{.}$
Our goal is to show that $\bar{\mu}_{u}>\hat{\mu}_{u}$ for some strictly-increasing, continuous, and concave $u$. Let us specify a particular such $u$. Let
$u(x)=\begin{cases}x,\quad&\text{if}\quad x\leq a_{\theta^{\prime}},\\ \iota x+(1-\iota)a_{\theta^{\prime}},\quad&\text{if}\quad x>a_{\theta^{\prime}}\text{.}\end{cases}$
for some $\iota\in(0,1)$, so $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)}{u\left(a_{\theta}\right)-u\left(b_{\theta}\right)+\iota\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)}-\frac{u\left(\hat{a}_{\theta}\right)-u\left(b_{\theta}\right)}{u\left(\hat{a}_{\theta}\right)-u\left(b_{\theta}\right)+\iota\left(b_{\theta^{\prime}}-a_{\theta^{\prime}}\right)+a_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}\text{.}$
If $b_{\theta}\geq a_{\theta^{\prime}}$, this equals
$\frac{a_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}+b_{\theta^{\prime}}-a_{\theta^{\prime}}}-0>0\text{,}$
when $\iota=0$. If $b_{\theta}<a_{\theta^{\prime}}$, this equals
$1-\frac{u\left(\hat{a}_{\theta}\right)-b_{\theta}}{u\left(\hat{a}_{\theta}\right)-b_{\theta}+a_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}>0\text{,}$
when $\iota=0$.∎
A.2 Theorem 3.6 Proof
Proof.
[…]
$b_{\theta}<\hat{a}_{\theta}<a_{\theta},\quad\text{and}\quad a_{\theta^{\prime}}<\hat{a}_{\theta^{\prime}}<b_{\theta^{\prime}}\text{.}$
Our goal is to show that $\bar{\mu}_{u}>\hat{\mu}_{u}$ for some strictly-increasing, continuous $u$. Take
$u(x)=\begin{cases}\iota x+(1-\iota)\hat{a}_{0},\quad&\text{if}\quad x\leq\hat{a}_{\theta},\\ x,\quad&\text{if}\quad x>\hat{a}_{\theta}\text{,}\end{cases}$
for some $\iota\in(0,1)$. Thus, $\bar{\mu}_{u}-\hat{\mu}_{u}$ equals
$\frac{a_{\theta}-\hat{a}_{\theta}-\iota\left(b_{\theta}-\hat{a}_{\theta}\right)}{a_{\theta}-\hat{a}_{\theta}-\iota\left(b_{\theta}-\hat{a}_{\theta}\right)+u(b_{\theta^{\prime}})-u(a_{\theta^{\prime}})}-\frac{\iota\left(\hat{a}_{\theta}-b_{\theta}\right)}{\iota\left(\hat{a}_{\theta}-b_{\theta}\right)+u(b_{\theta^{\prime}})-u(\hat{a}_{\theta^{\prime}})}\text{.}$
If $b_{\theta^{\prime}}\leq\hat{a}_{\theta}$, this equals
$1-\frac{\hat{a}_{\theta}-b_{\theta}}{\hat{a}_{\theta}-b_{\theta}+b_{\theta^{\prime}}-\hat{a}_{\theta^{\prime}}}>0\text{,}$
when $\iota=0$. If $b_{\theta^{\prime}}>\hat{a}_{\theta}$, it equals
$\frac{a_{\theta}-\hat{a}_{\theta}}{a_{\theta}-\hat{a}_{\theta}+b_{\theta^{\prime}}-u(a_{\theta^{\prime}})}-0>0\text{,}$
when $\iota=0$. ∎
A.3 Proposition 4.7 Proof
Proof.
Fix $a$ and recall the notation
$\begin{split}\mathcal{A}_{B}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>\max_{b\in B}b_{\theta}\right\},\quad\mathcal{B}_{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<\max_{b\in B}b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}_{B}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=\max_{b\in B}b_{\theta}\right\}\text{.}\end{split}$
We also introduce the following notation: for a fixed $b$,
$\begin{split}\mathcal{A}_{b}\coloneqq&\left\{\theta\in\Theta\colon a_{\theta}>b_{\theta}\right\},\quad\mathcal{B}_{b}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}<b_{\theta}\right\},\\ &\text{and}\quad\mathcal{C}_{b}\coloneqq\left\{\theta\in\Theta\colon a_{\theta}=b_{\theta}\right\}\text{.}\end{split}$
Evidently, $\mathcal{A}_{B}\subseteq\mathcal{A}_{b}$, $\mathcal{B}_{b}\subseteq\mathcal{B}_{B}$, and $\mathcal{C}_{b}\cap\mathcal{A}_{B}=\emptyset$.
[…]
It remains to show that $a\succ\max_{b\in B}$Rightarrow $\hat{a}\succ\max_{b\in B}$. The region of beliefs on which $\mathbb{E}_{\mu}u(a_{\theta})\geq\max_{b\in B}\mathbb{E}_{\mu}u(b_{\theta})$ is a polyhedral subset of $\Delta$, $P_{a}$. By the Krein-Milman theorem, any $\mu\in P_{a}$ can be written as a convex combination of the extreme points of $P_{a}$. Moreover, the only extreme points of $P_{a}$ for which $\mathbb{E}_{\mu}u(a_{\theta})>\max_{b\in B}\mathbb{E}_{\mu}u(b_{\theta})$ are the vertices of the simplex corresponding to degenerate distributions on $\theta$s in $\mathcal{A}_{B}$. Accordingly, any $\mu\in P_{a}$ for which $\mathbb{E}_{\mu}u(a_{\theta})>\max_{b\in B}\mathbb{E}_{\mu}u(b_{\theta})$ can be written as a convex combination of extreme points of $P_{a}$ with strictly positive weight on a set of such vertices. But then, for any such $\mu$, it must also be that $\mathbb{E}_{\mu}u(\hat{a}_{\theta})>\max_{b\in B}\mathbb{E}_{\mu}u(b_{\theta})$, as $\hat{a}>\max_{b\in B}b_{\theta}$ for all $\theta\in\mathcal{A}_{B}$ plus the already-shown fact that $\mathbb{E}_{\mu}u(\hat{a}_{\theta})\geq\max_{b\in B}\mathbb{E}_{\mu}u(b_{\theta})$ for all $\mu\in P_{a}$.
[…]
First, $\hat{a}>\max_{b\in B}b_{\theta}$ for all $\theta\in\mathcal{A}_{B}$ is clearly necessary, as otherwise we could just take the degenerate distribution on an offending $\theta\in\mathcal{A}_{B}$. Now let this hold but suppose for the sake of contraposition that there exists some $b\in B$ such that $\hat{a}$ does not dominate a mixture of $a$ and $b$.
For any pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{B}\times\mathcal{B}_{B}$, we say that action $b^{*}\in B$ is the boundary action if there exists $\bar{\gamma}\in\left(0,1\right)$ such that for all $\gamma<\bar{\gamma}$,
$(1-\gamma)a_{\theta}+\gamma a_{\theta^{\prime}}>\max_{b\in B}\left\{(1-\gamma)b_{\theta}+\gamma b_{\theta^{\prime}}\right\}$
and
$(1-\bar{\gamma})a_{\theta}+\bar{\gamma}a_{\theta^{\prime}}=(1-\bar{\gamma})b^{*}_{\theta}+\bar{\gamma}b^{*}_{\theta^{\prime}}\text{.}$
Claim A.1.
It is without loss of generality to assume that $b^{*}$ is unique. Moreover, for any concave and strictly increasing $u$ and all $\gamma\in\left[0,1\right]$,
$(1-\gamma)u(a_{\theta})+\gamma u(a_{\theta^{\prime}})\underset{(>)}{\geq}(1-\gamma)u(b^{*}_{\theta})+\gamma u(b^{*}_{\theta^{\prime}})\ \Rightarrow\ (1-\gamma)u(a_{\theta})+\gamma u(a_{\theta^{\prime}})\underset{(>)}{\geq}(1-\gamma)u(b_{\theta})+\gamma u(b_{\theta^{\prime}})\text{,}$
for all $b\in B$.
Proof.
Our assumption of single-peakedness means that on any $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{B}\times\mathcal{B}_{B}$, the only dominated actions for a risk-neutral agent are those that are dominated (on these two states) by some other action. Accordingly, the set of undominated actions on this pair of states does not grow in the DM’s risk aversion. On the other hand, the two Proposition 1s in Battigalli, Cerreia-Vioglio, Maccheroni, and Marinacci (2016) and Weinstein (2016) reveal that increased risk aversion can only increase (in a set-inclusion sense) the rationalizable set. Thus, the set of undominated actions on this pair of states is the same regardless of the DM’s concave $u$. Finally, we set $b^{*}$ to be the action with the largest value in state $\theta$ of the undominated actions (on these two states) in $B$, making an arbitrary selection if multiple such actions have the same payoffs in $\theta$ and $\theta^{\prime}$. ∎
This claim highlights the function of our assumption of single-peakedness. It ensures that regardless of the DM’s risk-averse utility, for each pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{B}\times\mathcal{B}_{B}$, there is only one other action that matters and it is the same regardless of $u$.
By Lemma 3.2, it suffices to show that there exists some pair $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{B}\times\mathcal{B}_{B}$ for which there does not exist a $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b^{*}_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b^{*}_{\theta^{\prime}}\text{,}$
where $b^{*}$ is the boundary action. We call this a boundary violation.
Claim A.2.
If $\hat{a}_{\theta^{\prime}}<a_{\theta^{\prime}}$ for some $\theta^{\prime}\in\mathcal{B}_{B}$, there exists a boundary violation.
Proof.
This is an immediate consequence of the fact that $b_{\theta^{\prime}}^{*}>a_{\theta^{\prime}}>\hat{a}_{\theta^{\prime}}$, so no averaging of the first two can produce something greater than the third. ∎
Henceforth, we assume that for all $\theta^{\prime}\in\mathcal{B}_{B}$, $\hat{a}_{\theta^{\prime}}\geq a_{\theta^{\prime}}$.
Claim A.3.
If there exists some $b\in B$ and $\theta^{\dagger}\in\mathcal{C}_{b}$ such that $\hat{a}_{\theta^{\dagger}}<a_{\theta^{\dagger}}$ then $\hat{a}$ is not $B$-superior to $a$.
Proof.
Suppose there exists some $b\in B$ and $\theta^{\dagger}\in\mathcal{C}_{b}$ such that $\hat{a}_{\theta^{\dagger}}<a_{\theta^{\dagger}}$. If $\hat{a}_{\theta}<a_{\theta}$ for some $\theta\in\mathcal{C}_{B}$, we have the desired conclusion. If $\hat{a}_{\theta}\geq a_{\theta}$ for all $\theta\in\mathcal{C}_{B}$ but there exists some $b\in B$ and $\theta\in\mathcal{C}_{b}$ such that $\hat{a}_{\theta}<a_{\theta}$, it must be the case that $\hat{a}_{\theta^{\prime}}<a_{\theta^{\prime}}$ for some $\theta^{\prime}\in\mathcal{B}_{B}$, a contradiction. ∎
As a result of Claim A.3, we assume that for all $\theta\in\mathcal{C}_{b}$, $\hat{a}_{\theta}\geq a_{\theta}$.
Claim A.4.
If there exists some $b\in B$ and $\theta\in\mathcal{A}_{b}$ such that $\hat{a}_{\theta}<b_{\theta}$, then $\hat{a}$ is not $B$-superior to $a$.
Proof.
Suppose there exists some $b\in B$ and $\theta\in\mathcal{A}_{b}$ such that $\hat{a}_{\theta}<b_{\theta}$. But then either $\theta\in\mathcal{A}_{B}$, $\theta\in\mathcal{C}_{B}$, or $\theta\in\mathcal{B}_{B}$ and $\hat{a}_{\theta}<b_{\theta}<a_{\theta}$; all contradictions.∎
As a result of Claim A.4, we assume that for all $\theta\in\mathcal{A}_{b}$, $\hat{a}_{\theta}\geq b_{\theta}$.
By Lemma 3.3, $\hat{a}$ does not pairwise dominate a collection of mixtures of $a$ and $b$. That is, on some $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{b}\times\mathcal{B}_{b}$ there does not exist a $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta^{\prime}}\text{.}$
As $\hat{a}_{\theta^{\prime}}\geq a_{\theta^{\prime}}$ for all $\theta^{\prime}\in\mathcal{B}_{B}$ this implies that on some $\left(\theta,\theta^{\prime}\right)\in\mathcal{A}_{B}\times\mathcal{B}_{B}$, there does not exist a $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b_{\theta^{\prime}}\text{.}$
We need to show that such a convex weight does not exist on this pair of states with respect to the boundary action $b^{*}$. But this is obvious: given $\hat{a}_{\theta^{\prime}}\geq a_{\theta^{\prime}}$, the non-existence of the weight $\lambda_{\theta,\theta^{\prime}}$ is equivalent to the line (in $\gamma$)
$(1-\gamma)\hat{a}_{\theta}+\gamma\hat{a}_{\theta^{\prime}}$
lying strictly below the line
$(1-\gamma)a_{\theta}+\gamma a_{\theta^{\prime}}\text{,}$
for all
$\gamma\leq\frac{a_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}+b_{\theta^{\prime}}-a_{\theta^{\prime}}}\text{,}$
which is the weight equalizing $a$ and $b$ on these two states. Moreover, by the definition of the boundary action, the weight equalizing $a$ and $b$ on these two states
$\frac{a_{\theta}-b^{*}_{\theta}}{a_{\theta}-b^{*}_{\theta}+b^{*}_{\theta^{\prime}}-a_{\theta^{\prime}}}\leq\frac{a_{\theta}-b_{\theta}}{a_{\theta}-b_{\theta}+b_{\theta^{\prime}}-a_{\theta^{\prime}}}\text{,}$
implying that there does not exist $\lambda_{\theta,\theta^{\prime}}\in\left[0,1\right]$ such that
$\hat{a}_{\theta}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b^{*}_{\theta}\quad\text{and}\quad\hat{a}_{\theta^{\prime}}\geq\lambda_{\theta,\theta^{\prime}}a_{\theta^{\prime}}+\left(1-\lambda_{\theta,\theta^{\prime}}\right)b^{*}_{\theta^{\prime}}\text{.}$
We have produced the desired violation, so we are done.∎
A.4 Proposition 5.2 Proof
Proof.
[…]
Suppose first that $\mu_{0}$ and $c$ are such that $p=1$. This is equivalent to $\mu_{0}\leq\bar{\mu}_{u}$ and
$\left(\alpha_{1}-\alpha_{0}-c^{\prime}(\mu_{0})\right)\mu+c^{\prime}(\mu_{0})\mu_{0}+\alpha_{0}\geq 0\text{,}$
for all $\mu\in\left[0,1\right]$. For any $\mu\in\left[0,1\right]$ this expression is strictly increasing in both $\alpha_{0}$ and $\alpha_{1}$, so we have
$\left(\hat{\alpha}_{1}-\hat{\alpha}_{0}-c^{\prime}(\mu_{0})\right)\mu+c^{\prime}(\mu_{0})\mu_{0}+\hat{\alpha}_{0}\geq 0\text{,}$
for all $\mu\in\left[0,1\right]$, i.e., $\hat{p}=1$.
Now suppose that $\mu_{0}$ is such that $p\in\left(0,1\right)$. In this case, the support of $F^{*}$, $\mu_{L}$ and $\mu_{H}$, solves
$\alpha_{1}-\alpha_{0}-c^{\prime}\left(\mu_{L}\right)+c^{\prime}\left(\mu_{H}\right)=0\text{,}$
and
$c^{\prime}\left(\mu_{L}\right)\mu_{L}-c^{\prime}\left(\mu_{H}\right)\mu_{H}+c\left(\mu_{H}\right)+\alpha_{0}-c\left(\mu_{L}\right)=0\text{.}$
Appealing to the implicit function theorem, we obtain
$\mu_{L}^{\prime}(\alpha_{1})=\frac{\mu_{H}}{c^{\prime\prime}\left(\mu_{L}\right)\left(\mu_{H}-\mu_{L}\right)}>0,\quad\text{and}\quad\mu_{H}^{\prime}(\alpha_{1})=\frac{\mu_{L}}{c^{\prime\prime}\left(\mu_{H}\right)\left(\mu_{H}-\mu_{L}\right)}>0\text{.}$
Likewise,
$\mu_{L}^{\prime}(\alpha_{1})=\frac{1-\mu_{H}}{c^{\prime\prime}\left(\mu_{L}\right)\left(\mu_{H}-\mu_{L}\right)}>0,\quad\text{and}\quad\mu_{H}^{\prime}(\alpha_{1})=\frac{1-\mu_{L}}{c^{\prime\prime}\left(\mu_{H}\right)\left(\mu_{H}-\mu_{L}\right)}>0\text{.}$
Evidently, $p$ is strictly increasing in both $\mu_{H}$ and $\mu_{L}$, so $\hat{p}\geq p$.
[…]
Case iii is immediate. Let the DM be risk neutral: then there exists a $\mu^{\prime}\in\left(0,1\right)$, such that for any belief $\mu<\mu^{\prime}$, it is uniquely optimal for the DM to take action $a$ when her menu is $\left\{a,b\right\}$ and $b$ when her menu is $\left\{\hat{a},b\right\}$. Then, one need only pick a sufficiently convex $c$–that such a convex $c$ can always be found is an implication of Lemma A.1 in Whitmeyer (2023) [^9] This is not strictly true as Whitmeyer (2023) does not impose that the cost function is twice continuously differentiable, merely strictly convex. However, it is easy to extend that result to smooth functions: see, e.g. (2019) (https://mathoverflow.net/users/943/dmitri panov). Alternatively, one could remove the twice-continuously differentiable specification at the expense of not being able to appeal to the implicit function theorem in the sufficiency portion of the proof.–such that for prior $\mu_{0}$, no learning is uniquely optimal in both problems, in which case $p=1>0=\hat{p}$.
Case i is also easy. Observe that $\hat{a}$ is not $b$-superior to $a$, so there exists a strictly increasing concave $u$ for which $10<\hat{\mu}_{u}<\bar{\mu}_{u}<$ 1. We maintain the convention $\alpha_{\theta}\equiv u(a_{\theta})$ and $\hat{\alpha}_{\theta}\equiv u(\hat{a}_{\theta})$ for all $\theta\in\left\{0,1\right\}$ and also introduce the notation $\beta_{\theta}\equiv u(b_{\theta})$ for all $\theta\in\left\{0,1\right\}$.
We tweak the notation
$\ell_{a}=\mu\alpha_{1}+\left(1-\mu\right)\alpha_{0},\ \ell_{\hat{a}}=\mu\hat{\alpha}_{1}+\left(1-\mu\right)\hat{\alpha}_{0},\text{ and }\ell_{b}=\mu\beta_{1}+\left(1-\mu\right)\beta_{0}\text{,}$
and define $W(\mu)\coloneqq\max\left\{\ell_{a},\ell_{\hat{a}},\ell_{b}\right\}$. We let $\tilde{\mu}$ denote the intersection of $\ell_{a}$ and $\ell_{\hat{a}}$, and observe that $0<\tilde{\mu}<\bar{\mu}$; this holds because $\ell_{\hat{a}}$ has a steeper slope and a strictly larger $y$-intercept than $\ell_{a}$. Then, Lemma A.1 in Whitmeyer (2023) implies that for any triple $\mu_{1}\in\left(0,\tilde{\mu}\right)$, $\mu_{2}\in\left(\tilde{\mu},\bar{\mu}\right)$, and $\mu_{3}\in\left(\bar{\mu},1\right)$, there exists a UPS cost such that when the DM’s value function is $W$, any optimal learning has support on the three specified points. Accordingly, for such a cost function, when $\mu_{0}=\mu_{2}$, $p=1$ and $\hat{p}<1$.
Finally, case ii: the argument from the previous paragraph allows us to assume that $\hat{a}$ is $b$-superior to $a$, or else we are done. Consequently, $\bar{\mu}_{u}\leq\hat{\mu}_{u}$ for all permissible $u$. Fix such a $u$ and normalize payoffs so that $\beta_{\theta}=0$ for all $\theta\in\left\{0,1\right\}$ (this is without loss of generality, as $u$ has been fixed). Now take a line $f(\mu)\coloneqq-\gamma\mu+\delta$, where $0<\delta<\alpha_{0}-\hat{\alpha}_{0}$, $\gamma>\delta$, and
$\frac{\delta}{\gamma}>\hat{\mu}_{u}=\frac{\hat{\alpha}_{0}}{\hat{\alpha}_{0}-\hat{\alpha}_{1}}\text{.}$
Defining $T(\mu)\coloneqq\max\left\{\ell_{a},\max\left\{\ell_{\hat{a}},0\right\}+f,0\right\}$, we note that this piecewise-affine curve has three kink points. First, at some $\mu_{1}\in\left(0,1\right)$, where $\ell_{a}$ and $\ell_{\hat{a}}+f$ intersect. Second, at some $\mu_{2}\in\left(\mu_{1},1\right)$, where 00 and $\ell_{\hat{a}}$ intersect. Third, at some $\mu_{3}\in\left(\mu_{2},1\right)$ where $f$ and 00 intersect. Again appealing to Lemma A.1 in Whitmeyer (2023), taking a prior $\mu_{0}\in\left(\mu_{2},\mu_{3}\right)$ we note the existence of a UPS cost producing $p>0$ and $\hat{p}=0$.∎